### v1

# Poincaré inequality on the monotone sets revisited

Abstract
In this short note, we provide a simple inductive proof of the Poincaré inequality on the monotone sets, which is firstly established by Fei and Pinto Jr. [2] using directed heat semigroup method. Such inequality improves the mixing time upper bound of the censored random walk on constant-density monotone sets from $O(n^{3})$ by Ding and Mossel [1] to $O(n^{2})$.
1 Introduction
Discrete-time random walks on finite graphs $G=(V,E)$ model stochastic processes where a walker iteratively transitions between adjacent vertices. Formally, at each timestep $t$, the walker moves from current vertex $u$ to neighbor $v$ with probability $P(u,v)=\frac{1}{{\deg}(u)}$, defining a Markov chain with state space $V$ and transition matrix $P$. The stationary distribution $\pi$, satisfying $\pi P=\pi$, is given by $\pi(v)=\frac{\deg(v)}{2|E|}$ for connected non-bipartite graph. The mixing time $t_{{\rm mix}}(\varepsilon)$ quantifies the convergence rate to stationarity, defined as the minimal $t$ such that:
$\max_{u\in V}\|P^{t}(u,\cdot)-\pi\|_{{\rm TV}}\leq\varepsilon,$ (1)
where $\|\cdot\|_{{\rm TV}}$ denotes the total variation, and $\varepsilon\in(0,1)$ governs approximation accuracy.
An important example is the $n$-dimensional discrete cube. This is a graph $Q_{n}=(\{0,1\}^{n},E)$ with $2^{n}$ vertices indexed by $n$-dimensional Boolean vector. Two 0-1 vectors are adjacent if they differ in exactly one coordinate. We say a set $A\subseteq\{0,1\}^{n}$ is monotone if $x\in A$ and $x\leq y$ imply that $y\in A$. Here, $x\leq y$ denotes $x_{i}\leq y_{i}$ for all $i\in[n]$. Ding and Mossel [1] introduced the study of the random walk on $\{0,1\}^{n}$ censored to $A$. This walk, when at state $x\in A$, attempts to flip a uniformly random coordinate; it moves to the neighbor if it resides in $A$, otherwise it stays at $x$. Note that if we do not add some properties on $A$, the censored random walk may mix well or extremely poorly even when $A$ is large and connected, see Example 1 and 2 in [2]. A significant open question (Question 1.1 in [1]) asked if the optimal mixing time $O_{c}(n\log n)$ holds for $\mu(A)\geq c$, where we use $\mu:=\mu_{1/2}^{\otimes n}$ to denote the uniform probability measure on the cube.
Question 1.1 (Question 1.1 in [1]).
Suppose that there exists a constant $c>0$ such that a monotone subset $A\subseteq\{0,1\}^{n}$ has measure $\mu(A)\geq c$. Consider the censored random walk on $A$ defined above. Is it true that $t_{{\rm mix}}(\varepsilon)\leq O_{\varepsilon,c}(n\log n)$?
When the density $\mu(A)$ is constant, Ding and Mossel established a mixing time bound of $O(n^{3})$. Very recently, Fei and Pinto Jr. [2] show an improved mixing time bound of $O(n^{2})$ for constant-density monotone sets via establishing the following Poincaré inequality on the non-product space.
Theorem 1.2 [2] Theorem 1.6).
Let $A\subseteq\{0,1\}^{n}$ be a non-empty monotone set. We have for all $f:A\to\mathbb{R}$,
${\rm Var}_{A}[f]\leq\frac{1}{1-\sqrt{1-\mu(A)}}\cdot\mathcal{E}_{A}(f).$ (2)
Here ${\rm Var}_{A}[f]$ stands for the variance of $f(x)$ where $x$ is a uniformly random element of $A$ and $\mathcal{E}_{A}(f)$ is called “restricted” Dirichlet form of a function on $A$ defined as follows
$\mathcal{E}_{A}(f)=\mathcal{E}_{A}(f,f):=\frac{1}{4}\cdot\sum\limits_{i=1}^{n}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right],$ (3)
where the expectation is taken with respect to the uniform probability measure on the monotone subset $A\subseteq\{0,1\}^{n}$ and $x^{\oplus i}$ is obtained from $x$ by flipping the $i$th coordinate.
Note that when $A=Q_{n}$, the above result recovers the classical Poincaré inequality on the hypercube. Their proof relies heavily on the connections between the directed heat flow and the directed Poincaré inequality. In this note, we provide a simple inductive proof of their Poincaré inequality, though yielding a slightly worse constant (theirs being 1 versus ours at 2). Nevertheless, we still obtain an $O(n^{2})$ mixing time bound for monotone sets $A$ of constant density via standard Markov chain theory (Theorem 12.4 in [3]). Specifically, we will prove the following theorem, a weaker version of Theorem 1.2.
Theorem 1.3.
Let $A\subseteq\{0,1\}^{n}$ be a non-empty monotone set. We have for all $f:A\to\mathbb{R}$,
${\rm Var}_{A}[f]\leq\frac{2}{1-\sqrt{1-\mu(A)}}\cdot\mathcal{E}_{A}(f).$ (4)
2 Preliminaries and Proofs
In this section, we outline some notation and terminology, and present some simple lemmas for inductive proof of Theorem 1.3.
Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V^{n}$. Let
$V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$
and
$A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},\quad A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}.$
Let $a_{0}:=\mu^{n-1}(A_{0}),a_{1}:=\mu^{n-1}(A_{1})$, and note that
$\begin{split}a&:=\mu^{n}(A)=\frac{1}{2^{n}}\sum_{x\in V^{n}}\mathbbm{1}_{A}(x)=\frac{1}{2}\cdot\frac{1}{2^{n-1}}\left(\sum_{(x^{\prime},1)\in V_{1}}\mathbbm{1}_{A}(x^{\prime},1)+\sum_{(x^{\prime},0)\in V_{0}}\mathbbm{1}_{A}(x^{\prime},0)\right)\\ &=\frac{1}{2}\cdot\frac{1}{2^{n-1}}\left(\sum_{x^{\prime}\in V^{n-1}}\mathbbm{1}_{A_{1}}(x^{\prime})+\sum_{x^{\prime}\in V^{n-1}}\mathbbm{1}_{A_{0}}(x^{\prime})\right)=\frac{1}{2}\mu^{n-1}(A_{1})+\frac{1}{2}\mu^{n-1}(A_{0})=\frac{a_{0}+a_{1}}{2}.\end{split}$ (5)
Define $f_{0}:A_{0}\to\mathbb{R},\ f_{1}:A_{1}\to\mathbb{R}$ by
$f_{0}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},0)\quad\text{and}\quad f_{1}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},1).$
Fact 2.1.
If $A$ is an increasing set, then $A_{0}\subset A_{1}$.
Proof.
For any $x^{\prime}\in A_{0}$, we shall see that $x^{\prime}\in A_{1}$. It is easy to see because $(x^{\prime},0)\leq(x^{\prime},1)$ and $A$ is increasing. ∎
By applying conditional expectation expansion to $n$th variable of restricted Dirichlet forms, we can readily reduce the Dirichlet form of the original function to that of the restricted functions.
Lemma 2.2.
For any $f:A\to\mathbb{R}$, we have
$\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$ (6)
Proof.
We use $x^{\prime}$ to denote the first $(n-1)$-coordinates of $x$. Note that
$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\sum\limits_{i=1}^{n-1}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right]+\frac{1}{4}\cdot\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus n}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1-x_{n})\in A\}}\right]\\ &=\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},1)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},0)\in A\}}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1)\in A\}}\right]\\ &=\frac{a_{1}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{1}\}}\right]+\frac{a_{0}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{0}\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right],\end{split}$
where in the second equality, we observe that
$\begin{split}\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}&\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\right]=\frac{1}{|A|}\cdot\sum\limits_{x^{\prime}\in A_{1}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\\ &=\frac{1}{|A|}\cdot\sum\limits_{x^{\prime}\in A_{0}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}=\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right].\end{split}$
Similarly, we can reduce the variance of the original function to that of the restricted functions.
Lemma 2.3.
For any $f:A\to\mathbb{R}$, we have
${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (7)
Proof.
Note that
$\begin{split}{\rm Var}_{A}[f]=\underset{x\sim A}{\mathbb{E}}[|f-\underset{x\sim A}{\mathbb{E}}f|^{2}]=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\end{split}$
Let $\mu=\underset{x\sim A}{\mathbb{E}}[f(x)],\mu_{1}=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]$ and $\mu_{0}=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]$. Then
$\begin{split}\mu&=\frac{1}{|A|}\sum_{x\in A}f(x)=\frac{1}{|A|}\sum_{x^{\prime}\in A_{1}}f(x^{\prime},1)+\frac{1}{|A|}\sum_{x^{\prime}\in A_{0}}f(x^{\prime},0)=\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]+\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\\ &=\frac{a_{1}}{2a}\mu_{1}+\frac{a_{0}}{2a}\mu_{0}.\end{split}$ (8)
Thus
$\begin{split}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]&=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\frac{a_{1}}{2a}\mu_{1}-\frac{a_{0}}{2a}\mu_{0}\right|^{2}\right]=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}+\mu_{1}-\frac{a_{1}}{2a}\mu_{1}-\frac{a_{0}}{2a}\mu_{0}\right|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}+\frac{a_{0}}{2a}(\mu_{1}-\mu_{0})\right|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}\right|^{2}\right]+2\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}-\mu_{1})\cdot\frac{a_{0}}{2a}(\mu_{1}-\mu_{0})\right]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}\\ &={\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}.\end{split}$
Similarly, we have
$\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]={\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}.$
Thus we have
$\begin{split}{\rm Var}_{A}[f]&=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{1}a_{0}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}^{2}+a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(\mu_{1}-\mu_{0})^{2}\end{split}$
3 An inductive proof of Theorem 1.3
Our goal is to prove that for all $f:A\to\mathbb{R}$,
$\mathcal{E}_{A}(f)\geq c\cdot(1-\sqrt{1-a})\cdot{\rm Var}_{A}[f],$ (9)
where $c>0$ is a universal constant (we can take $c=0.5$). It is easy to check that the theorem holds for $n=1$. Let $n\geq 2$, and suppose the statement of the theorem holds when $n$ is replaced by $n-1$. Using Lemma 2.2 to split the Dirichlet form of the original function to two restricted functions:
$\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$ (10)
By inductive assumption, we have
$\begin{split}&\mathcal{E}_{A_{1}}(f_{1})\geq c\cdot(1-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}],\\ &\mathcal{E}_{A_{0}}(f_{0})\geq c\cdot(1-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}].\end{split}$ (11)
Thus our goal is to show that
$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\geq\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}(c-c\sqrt{1-a}){\rm Var}_{A}[f].\end{split}$ (12)
We also use Lemma 2.3 to split the variance of the original function to two restricted functions:
${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}^{2}+a_{0}a_{1}^{2}}{8a^{3}}\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (13)
Thus it suffices to show that
$\begin{split}&\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(c-c\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (14)
It is equivalent to show that
$\begin{split}&ca_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+ca_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}c\frac{a_{1}a_{0}}{a_{0}+a_{1}}(1-\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (15)
Lemma 3.1.
We may assume both $f_{0},f_{1}$ to be constant on $A_{0}$ and on $A_{1}\setminus A_{0}$.
Proof.
Let $\mu_{1}:=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]$, $\alpha:=\mathbb{E}_{x^{\prime}\sim A_{0}}[f_{1}]$, $\beta:=\mathbb{E}_{x^{\prime}\sim A_{1}\setminus A_{0}}[f_{1}]$ and $\gamma:=\mathbb{E}_{x^{\prime}\sim A_{0}}[f_{0}]$. Note that
$\begin{split}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}f_{1}|^{2}]&=\frac{|A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}+\frac{|A_{1}\setminus A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{1}\setminus A_{0}|}\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}\\ &=\frac{a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}],\end{split}$ (16)
and
$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}(x^{\prime})-\mu_{0}-(f_{1}(x^{\prime})-\mu_{1})+\mu_{0}-\mu_{1}|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}-\mu_{0}|^{2}\right]+\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{1}-\mu_{1}|^{2}\right]+2\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]+(\mu_{0}-\mu_{1})^{2}+2(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (17)
For the sake of simplicity, we denote $C_{1}=\frac{1}{2}\cdot a_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}})$ and $C_{0}=\frac{1}{2}\cdot a_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}})$. Then
$\begin{split}&C_{1}\cdot{\rm Var}_{A_{1}}[f_{1}]+C_{0}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+C_{1}\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]\\ &+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]+\frac{a_{0}}{2}(\mu_{0}-\mu_{1})^{2}+a_{0}(\mu_{0}-\mu_{1})(\mu_{1}-\alpha)\end{split}$ (18)
Now we focus on the term
$\begin{split}&\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &=\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)|\mu_{1}-f_{1}(x^{\prime})|^{2}+\left(C_{0}+\frac{a_{0}}{2}\right)|f_{0}(x^{\prime})-\mu_{0}|^{2}+a_{0}(f_{0}(x^{\prime})-\mu_{0})(\mu_{1}-f_{1}(x^{\prime}))\\ &=\frac{1}{|A_{0}|}\sum_{x^{\prime}\in A_{0}}\vec{b}_{x^{\prime}}^{\top}G_{x^{\prime}}\vec{b}_{x^{\prime}},\end{split}$ (19)
where
$\vec{b}_{x^{\prime}}=\begin{bmatrix}f_{0}(x^{\prime})-\mu_{0}\\ \mu_{1}-f_{1}(x^{\prime})\end{bmatrix},\\ G_{x^{\prime}}=\begin{bmatrix}a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}&\frac{a_{0}}{2}\\ \frac{a_{0}}{2}&a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}.\end{bmatrix}\$
We claim that the matrix $G_{x^{\prime}}$ is positive semidefinite. Note that $G_{x^{\prime}}$ is a symmetric matrix, then we shall use the basic fact that a symmetric matrix $G=\begin{bmatrix}b&c\\ c&d\end{bmatrix}\in\mathbb{R}^{2\times 2}$ is positive semidefinite if and only if $b\geq 0,d\geq 0$ and $bd-c^{2}\geq 0$.
It is easy to see that $a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}\geq 0$ and $a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}\geq 0$ as $a_{0}\leq a\leq a_{1}$. Thus it suffices to verify
$\frac{a_{0}^{2}}{4}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-\frac{a_{0}^{2}}{4}\geq 0.$ (20)
Let $s:=\sqrt{1-a_{0}}$ and $t:=\sqrt{1-a_{1}}$, then $0\leq t\leq s\leq 1$, $21-a=(s^{2}+t^{2})/$ 2 and
$\begin{split}&(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-1\\ &=\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-s+1\right)\cdot\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-t+1\right)-1\\ &=-\frac{1}{2}\cdot(s+t-2)\cdot(\sqrt{2}\sqrt{s^{2}+t^{2}}-s-t)\geq 0,\end{split}$ (21)
where the final one is obviously nonnegative as $s+t-2\leq 0$ and $2(s^{2}+t^{2})\geq(s+t)^{2}$. By the second-order condition of convexity, we know that for any $x^{\prime}\in A_{0}$, the quadratic function $G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))$ is a convex function. Thus by Jensen’s inequality, we have
$\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))]\geq G\left(\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})-\mu_{0}],\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[\mu_{1}-f_{1}(x^{\prime})]\right)=G(0,\mu_{1}-\alpha),$
and
$\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]\geq\left(\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[f_{1}-\mu_{1}]\right)^{2}=(\beta-\mu_{1})^{2}.$
Hence replacing $f_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of (15). ∎
Now let $\alpha$ be the value of $f_{1}$ on $A_{0}$ and $\beta$ be the value of $f_{1}$ on $A_{1}\setminus A_{0}$. Let $\gamma$ be the value of $f_{0}$ on $A_{0}$. Then we compute
$\begin{split}&{\rm Var}_{A_{1}}[f_{1}]=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}^{2}]-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}]^{2}\\ &=\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta^{2})-\left(\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta)\right)^{2}\\ &=\frac{1}{a_{1}}(a_{0}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot\beta^{2})-\frac{1}{a_{1}^{2}}(a_{0}\cdot\alpha+(a_{1}-a_{0})\cdot\beta)^{2}\\ &=\frac{a_{1}a_{0}-a_{0}^{2}}{a_{1}^{2}}\cdot(\alpha-\beta)^{2},\end{split}$ (22)
and
${\rm Var}_{A_{0}}[f_{0}]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}^{2}]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}]^{2}=\frac{a_{0}\cdot 2^{n-1}\cdot\gamma^{2}}{2^{n-1}\cdot a_{0}}-\left(\frac{a_{0}\cdot 2^{n-1}\cdot\gamma}{2^{n-1}\cdot a_{0}}\right)^{2}=0.$ (23)
Thus our goal is to prove that
$\begin{split}&c(\sqrt{1-a}-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})\cdot(\alpha-\beta)^{2}+\frac{a_{1}}{2}\cdot(\gamma-\alpha)^{2}\\ &\geq\frac{c}{a_{1}+a_{0}}\cdot(1-\sqrt{1-a})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}.\end{split}$ (24)
Note that
$\sqrt{1-a}-\sqrt{1-a_{1}}=\frac{1-a-(1-a_{1})}{\sqrt{1-a}+\sqrt{1-a_{1}}}=\frac{a_{1}-\frac{a_{1}+a_{0}}{2}}{\sqrt{1-a}+\sqrt{1-a_{1}}}=\frac{a_{1}-a_{0}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})},$ (25)
and
$1-\sqrt{1-a}=\frac{1-(1-a)}{1+\sqrt{1-a}}=\frac{a_{1}+a_{0}}{2(1+\sqrt{1-a})}.$ (26)
Then it suffice to prove that
$\begin{split}&\frac{c\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})}+\frac{a_{1}}{2}\cdot(\alpha-\gamma)^{2}\geq\frac{c}{2(1+\sqrt{1-a})}\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\\ &\Leftrightarrow c\cdot(1+\sqrt{1-a})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\geq 0.\end{split}$ (27)
Note that
$a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)=(a_{0}-a_{1})\cdot(\alpha-\beta)+a_{1}\cdot(\alpha-\gamma)$
After some simple calculations, the following inequality we need to verify:
$\begin{split}&c\cdot(1-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\cdot(a_{0}-a_{1})\cdot(\alpha-\beta)\cdot(\alpha-\gamma)\geq 0.\end{split}$ (28)
If $\alpha=\gamma$, then it is trivially true. Thus we can assume that $\alpha\neq\gamma$ and let $T:=\frac{\alpha-\beta}{\alpha-\gamma}\cdot(a_{1}-a_{0})$. Then the above inequality transforms to a quadratic inequality in $T$:
$\begin{split}&c\cdot(1-\sqrt{1-a_{1}})\cdot T^{2}+2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot T\\ &a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\geq 0\end{split}$ (29)
Now let us rewrite the Left-hand side of the above inequality as a quadratic in $T$ with coefficients depending only on $a_{1},a_{0}$:
$\begin{split}&A:=c\cdot(1-\sqrt{1-a_{1}})\geq 0\\ &B:=2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\\ &C:=a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}}).\end{split}$
It only needs to verify $\Delta=B^{2}-4AC\leq 0$ with $c=0.5$, which will conclude the proof.
Lemma 3.2.
$B^{2}-4AC\leq 0$.
Proof.
Denote $x=a_{0},y=a_{1}$. Simplifying and rearranging, we need to prove that for any $0\leq x\leq y\leq 1$, we have
$\begin{split}&y\cdot\left(\sqrt{1-\frac{x+y}{2}}+\sqrt{1-y}\right)-2(1-\sqrt{1-y})\cdot\left(1+\sqrt{1-\frac{x+y}{2}}-\frac{y}{2}\right)\leq 0\\ &\Leftrightarrow\left(\sqrt{1-\frac{x+y}{2}}+1\right)\cdot(y-2+2\sqrt{1-y})\leq 0.\end{split}$ (30)
It suffices to show that $(y-2+2\sqrt{1-y})\leq 0$. It is trivially true as
$\begin{split}&y-2+2\sqrt{1-y}\leq 0\\ &\Leftrightarrow 2\sqrt{1-y}\leq 2-y\\ &\Leftrightarrow y^{2}\geq 0.\end{split}$ (31)

### v2

@@
-Poincaré inequality on the monotone sets revisited
+Functional inequalities on the biased and restricted cube: an induction approach
 Abstract
-In this short note, we provide a simple inductive proof of the Poincaré inequality on the monotone sets, which is firstly established by Fei and Pinto Jr.
-[2] using directed heat semigroup method.
-Such inequality improves the mixing time upper bound of the censored random walk on constant-density monotone sets from $O(n^{3})$ by Ding and Mossel [1] to $O(n^{2})$.
+We develop new discrete functional inequalities on the hypercube via the induction-by-restrictions method.
+This method reduces high-dimensional inequalities to explicit low-dimensional analytic verifications and has recently proved effective for many discrete functional inequalities.
+We establish two results along this framework.
+First, we prove a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions, which recovers the classic biased edge-isoperimetric inequality for increasing sets and identifies increasing subcubes as the extremizers.
+This result also admits a probabilistic interpretation in terms of maximizing the mean first exit time of biased random walks.
+Second, we give an inductive proof of a Poincaré inequality on increasing subsets of the cube that was recently established by Fei and Ferreira Pinto Jr [17], yielding an $O(n^{2})$ upper bound on the mixing time of censored random walks, improving upon previous bounds.
 1 Introduction
+Functional inequalities on the discrete cube, such as the Poincaré, log-Sobolev, and edge-isoperimetric inequalities, form a central part of modern discrete analysis.
+They connect Boolean function analysis, discrete isoperimetric problems, and the spectral theory of Markov chains, providing powerful tools for studying concentration of measure, threshold phenomena, and mixing times of random walks.
+In the classical uniform product setting on $\{0,1\}^{n}$, these inequalities are well understood: sharp constants are known (e.g., the exact log-Sobolev constant [6]) and elegant proofs arise via tensorization [20], semigroup methods [23, 24], or discrete Fourier analysis [13].
+However, once one departs from the product setting–by restricting to structured subsets or by working under biased measures–classical approaches often fail or lose sharpness.
+Recent progress highlights both the challenges and the opportunities in these non-product environments: FKN-type structure theorems and sharp junta thresholds for low-degree Boolean functions on the slice/Johnson graph [18], (near) sharp log-Sobolev inequalities on the multislice [19], and representation-theoretic analyses of low-degree functions on the symmetric group that yield asymptotically sharp edge-isoperimetric bounds for the transposition graph [14].
+These developments motivate the search for new discrete functional inequalities adapted to non-product environments.
+A case in point that illustrates both the promise and the limitations of classical tools is the following.
+On the uniform cube, the logarithmic Sobolev inequality is tight for general real-valued functions, yet when specialized to indicator functions it recovers the edge-isoperimetric bound only up to a suboptimal multiplicative constant (by a factor $1/\ln 2$).
+Samorodnitsky [30] overcame this mismatch by proving a discrete functional inequality tailored to subsets: for functions supported on $A\subseteq\{0,1\}^{n}$ it yields the sharp Harper’s constant and characterizes equality via codimension-$k$ subcubes.
+Although Samorodnitsky’s inequality is formulated on the hypercube, recovering Harper’s optimal constant forces the right–hand side to involve the conditional average over the support set $A$.
+This quantity depends on the global geometry/size of $A$ and is not additive over coordinates; it cannot be written as an average of one-coordinate functionals.
+Consequently, the standard tensorization machinery for product spaces, which relies on coordinatewise factorization, cannot “lock onto” the correct constant.
+In this sense, even though the ambient space is product, conditioning on an arbitrary $A$ unfortunately destroys tensor structure at the level of the inequality, revealing difficulties that mirror those in genuinely non-product settings.
+In the $p$-biased setting, while a sharp edge-isoperimetric inequality for increasing sets is known (see, e.g., [15, 25]), the natural $p$-biased log-Sobolev inequalities (e.g., [7, 9]) again fail to deliver the optimal constant for indicators except in limiting regimes (such as $p\to 0$), leaving a quantitative and conceptual gap.
+These considerations highlight an urgent need to establish a Samorodnitsky-type functional inequality on the $p$-biased cube.
+A second illustration comes from mixing of censored random walks, where product structure is explicitly broken.
+If we censor the simple random walk on $\{0,1\}^{n}$ to a subset $A$ (only moves that stay in $A$ are accepted), the chain now lives on a non-product space with geometry dictated by the boundary of $A$.
+In this setting, standard product-based tools (tensorization, Fourier decompositions) no longer apply cleanly, and the behavior can vary widely: the walk may mix well, but it can also mix very poorly even when $A$ is large and connected (see [17, Examples 1–2]).
+Motivated by the question of Ding and Mossel [10, Question 1.1]—whether $t_{\mathrm{mix}}(\varepsilon)\leq O_{c,\varepsilon}(n\log n)$ holds for increasing sets of constant density—recent progress of Fei and Ferreira Pinto Jr [17] proves an $O(n^{2})$ bound by establishing a Poincaré inequality on the restricted (non-product) space.
+Their argument relies on a rather complex directed heat-flow method.
+This naturally raises a question: can simpler methods–such as an induction argument on the dimension–still be effective in such non-product spaces?
+Pursuing this direction can clarify which obstacles are intrinsic to the loss of product structure and offer a cleaner pathway toward functional inequalities in genuinely non-product settings.
+A particularly powerful paradigm that has recently emerged for proving such inequalities is the induction by restrictions combined with two-point inequalities, an approach widely adopted within the combinatorics and discrete analysis community, as exemplified by [1, 2, 3, 8, 12, 15, 16, 25, 30, 33].
+Our contributions.
+In this paper we establish two new discrete functional inequalities along this framework: one concerns a sharp $p$-biased edge-isoperimetric inequality for all real-valued increasing functions, and the other is a Poincaré inequality on increasing subsets of the hypercube.
+Both are proved by induction on the dimension combined with a two-point inequality, illustrating the versatility of this method across departures from the product setting.
+  • (1)
+    $p$-biased edge-isoperimetric inequality.
+In the uniform cube, Harper’s edge-isoperimetric theorem identifies subcubes as the unique minimizers of edge boundary, and Samorodnitsky [30] established a functional extension with sharp constants.
+For the $p$-biased hypercube endowed with product measure $\mu_{p}$, however, only weaker forms were known, typically losing sharp constants.
+We establish a new $p$-biased edge-isoperimetric inequality for real-valued increasing functions via induction on dimension and a two-point verification.
+Our result recovers the sharp $p$-biased isoperimetric inequality for increasing sets, and admits a probabilistic interpretation: increasing subcubes maximize the mean first exit time of the $p$-biased random walk among increasing sets of fixed measure.
+  • (2)
+    Poincaré inequality on increasing sets.
+Ding and Mossel [10] introduced the censored random walk on an increasing subset $A\subseteq\{0,1\}^{n}$ and asked whether it mixes in $O(n\log n)$ steps under constant density.
+They proved an $O(n^{3})$ upper bound.
+Fei and Ferreira Pinto Jr [17] established a Poincaré inequality on increasing sets that improves the spectral gap and implies an $O(n^{2})$ mixing-time upper bound.
+We give a simpler inductive proof of this Poincaré inequality, again leading to the $O(n^{2})$ mixing bound for the censored random walk.
+1.1 Functional version of the biased edge-isoperimetric inequality
+In this subsection we work on the $n$-dimensional hypercube.
+The state space is the graph $Q^{n}=(\{0,1\}^{n},E)$ with $2^{n}$ vertices indexed by binary strings, where two vertices are adjacent (written $x\sim y$) iff they differ in exactly one coordinate.
+Let $g:\{0,1\}^{n}\to\mathbb{R}$, we define the Dirichlet form of $g$ as
+$\mathcal{E}(g,g)=\underset{x\sim\mu}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}(g(x)-g(y))^{2}\right],$
+where the expectation is taken with respect to the uniform probability measure $\mu:=\mu_{1/2}^{\otimes n}$ on the hypercube.
+Furthermore, we denote by $x\sim A$ that $x$ is drawn uniformly at random from the subset $A\subseteq\{0,1\}^{n}$.
+In 2017, Samorodnitsky [30] established the following discrete functional inequality on the hypercube.
+Theorem 1.1 (Samorodnitsky [30]).
+Let $A$ be a subset of $\{0,1\}^{n}$ and let $g:\{0,1\}^{n}\to\mathbb{R}$ supported on $A$. Then
+$\mathcal{E}(g,g)\geq\frac{|A|}{2^{n-1}}\left(\underset{x\sim A}{\mathbb{E}}[|g(x)|]\right)^{2}\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right),$ (1)
+equality holds if and only if $g$ is an indicator function of a co-dimension $k$ subcube for any $0\leq k\leq n$.
+A subcube of co-dimension $k$ is a set of the form
+$\{x\in\{0,1\}^{n}:x_{i_{1}}=a_{1},x_{i_{2}}=a_{2},\dots,x_{i_{k}}=a_{k}\},$
+where $i_{1}<\cdots<i_{k}$ are coordinate indices, and $a_{1},\dots,a_{k}$ are fixed elements of $\{0,1\}$.
+The coordinates indexed by $i_{1},\dots,i_{k}$ are called the fixed coordinates; the other coordinate indices are called the active coordinates.
+Remark.
+  • (1)
+    When $g$ is an indicator function of $A$, the Samorodnitsky’s inequality (1) implies the classic edge-isoperimetric inequality on $Q^{n}$: for every $A\subset\{0,1\}^{n}$,
+    $|\nabla A|\geq|A|\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right),$ (2)
+    where the edge boundary $\nabla(A)$ is the set of edges between $A$ and its complement $A^{c}=V(Q^{n})\setminus A$.
+Equality in (2) holds if and only if $A$ is a subcube of co-dimension $k$, for any $0\leq k\leq n$.
+It is easy to see that $\mathcal{E}(\mathbbm{1}_{A},\mathbbm{1}_{A})=\frac{2|\nabla A|}{2^{n}}$.
+The fully edge isoperimetric problem for $Q^{n}$ was solved by Harper [21], Lindsey [28], Bernstein [5], and Hart [22], which states that among all subsets of $\{0,1\}^{n}$ of size $1\leq k\leq 2^{n}$, the first $k$ elements of the lexicographic ordering on $\mathcal{P}([n])$ (defined by $x<y\Leftrightarrow\max(x\Delta y)\in y$) have the smallest possible edge-boundary.
+  • (2)
+    A primary motivation for Samorodnitsky’s functional inequality is to obtain a general discrete functional inequality for real-valued functions that recovers the sharp/correct constant in (2).
+In contrast, the classical logarithmic Sobolev inequality on the hypercube only recovers (2) up to a multiplicative factor $\frac{1}{\ln 2}$:
+    $\mathcal{E}(g,g)\geq 2\cdot{\rm Ent}(g^{2}):=2\mathbb{E}[g^{2}\ln(g^{2})]-2\mathbb{E}[g^{2}]\ln(\mathbb{E}[g^{2}]).$ (3)
+  • (3)
+    To see the connection between Samorodnitsky’s inequality (1) and the log-Sobolev inequality (3), we rewrite it as
+    $\mathcal{E}(g^{\prime},g^{\prime})\geq 2\cdot\frac{|A|}{2^{n}}\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right)$
+    for nonnegative $g$, where $g^{\prime}:=g/\mathbb{E}_{A}[g]$, and $\mathbb{E}_{A}[g]$ is the conditional expectation on $A$.
+We generalize Samorodnitsky’s inequality to the $p$-biased setting so that the resulting functional inequality recovers the sharp constant in the $p$-biased edge-isoperimetric inequality.
+Let $f\in L^{2}(\{0,1\}^{n},\mu_{p}^{\otimes n})$, where $\mu_{p}$ is the Bernoulli measure on $\{0,1\}$ with $\mu_{p}(1)=p$ and $\mu_{p}(0)=1-p$.
+For $x\in\{0,1\}^{n}$, write $|x|=|\{i\in[n]:x_{i}=1\}|$, so that
+$\mu_{p}(x):=\mu_{p}^{\otimes n}(x)=p^{|x|}(1-p)^{n-|x|},\quad\mathbb{E}_{p}[f]:=\underset{x\sim\mu_{p}}{\mathbb{E}}[f(x)]=\sum\limits_{x\in\{0,1\}^{n}}f(x)\mu_{p}(x).$
+Many applications (e.g., to the study of percolation [4], threshold phenomena in random graphs [26, 31], and hardness of approximation [11]) rely upon the use of the biased measure on the discrete cube.
+The definition of Dirichlet quadratic form of $g$ with respect to the biased measure is
+$\mathcal{E}_{p}(g,g):=\underset{x\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}(g(x)-g(y))^{2}\right].$
+We consider increasing functions defined on the discrete cube $\{0,1\}^{n}$ endowed with the $p$-biased measure.
+Definition 1.2.
+A function $g:(\{0,1\}^{n},\mu_{p})\to\mathbb{R}$ is increasing if for all $x,y\in\{0,1\}^{n}$,
+$x_{i}\leq y_{i},\ \forall i\in[n]\Rightarrow g(x)\leq g(y).$
+A subset $A\subset\{0,1\}^{n}$ is called increasing if its characteristic function is increasing.
+The following ‘biased’ generalization of (2) for increasing subsets is considered folklore (see [15, 25]).
+Theorem 1.3.
+Let $A$ be an increasing subset of $\{0,1\}^{n}$ and let $0<p<1$. Then
+$p\cdot\mathcal{E}_{p}(\mathbbm{1}_{A},\mathbbm{1}_{A})\geq\mu_{p}(A)\log_{p}\mu_{p}(A),$ (4)
+equality holds in (4) if and only if $A$ is an increasing subcube.
+Inspired by (1), it is natural to seek a functional inequality on the $p$-biased cube for general real-valued (increasing) functions that, in the special case of indicator functions, reduces to the $p$-biased edge-isoperimetric inequality with the sharp constant.
+We establishes such an inequality, which recovers (4) by substituting $g=\mathbbm{1}_{A}$.
+Theorem 1.4.
+Let $A$ be a subset of $\{0,1\}^{n}$ and let $g$ be a real-valued increasing function on $\{0,1\}^{n}$ supported on $A$. Then for any $0<p<1$,
+$p\cdot\mathcal{E}_{p}(g,g)\geq\frac{\mathbb{E}_{p}[|g|]^{2}}{\mu_{p}(A)}\log_{p}\mu_{p}(A),$ (5)
+equality holds if and only if $g$ is an indicator function of an increasing subcube of co-dimension $k$ for any $0\leq k\leq n$.
+Remark.
+Note that the logarithmic Sobolev inequality (3) can be generalized to $p$-biased measure as follows, see [7, 9]:
+$\mathcal{E}_{p}(f,f)\geq\frac{1}{p(1-p)}\cdot\frac{1-2p}{\ln(1-p)-\ln(p)}\cdot{\rm Ent}_{p}(f^{2}).$ (6)
+When $f:=\mathbbm{1}_{A}$, we have the following
+$p\cdot\mathcal{E}_{p}(f,f)\geq\frac{\ln(1/p)}{1-p}\cdot\frac{1-2p}{\ln(1-p)-\ln(p)}\cdot\mu_{p}(f)\log_{p}\mu_{p}(f),$ (7)
+and hence for $p\to 0$,
+$(1+o_{p}(1))p\cdot\mathcal{E}_{p}(f,f)\geq\mu_{p}(f)\log_{p}\mu_{p}(f),$
+which again can not cover the original $p$-biased version of edge-isoperimetric inequality (4).
+We view the $p$-biased hypercube as a weighted undirected graph $Q^{n}_{p}=(\{0,1\}^{n},E,w_{p})$, where for every vertex $x\in\{0,1\}^{n}$ and coordinate index $i\in[n]$, the edge weight is defined by
+$\begin{split}w_{p}(\{x,x\oplus e_{i}\})&=p^{|x|}(1-p)^{n-|x|}+p^{|x\oplus e_{i}|}(1-p)^{n-|x\oplus e_{i}|}\\ &=p^{\sum_{j\in[n]:j\neq i}x_{j}}(1-p)^{n-\sum_{j\in[n]:j\neq i}x_{j}}\left(p^{x_{i}}(1-p)^{-x_{i}}+p^{x_{i}\oplus e_{i}}(1-p)^{-x_{i}\oplus e_{i}}\right)\\ &=p^{\sum_{j\in[n]:j\neq i}x_{j}}(1-p)^{n-\sum_{j\in[n]:j\neq i}x_{j}-1}.\end{split}$
+Thus each edge is weighted proportionally to the combined $p$-biased mass of its two endpoints.
+A random walk on a graph is a process that begins at some vertex, and at each time step moves to another vertex.
+When the graph is unweighted, the vertex the (simple) random walk moves to is chosen uniformly at random among the neighbors of the present vertex.
+When the graph is weighted, it moves to a neighbor with probability proportional to the weight of the corresponding edge.
+In [30], Samorodnitsky observed that the functional inequality (1) is equivalent to a probabilistic statement about random walks on the hypercube: subcubes maximize mean first exit time among all subsets of the same cardinality.
+Using the same argument, our new $p$-biased edge-isoperimetric inequality (5) admits an analogous probabilistic interpretation in terms of a simply random walk on the weighted hypercube $Q^{n}_{p}$.
+Let $A\subseteq\{0,1\}^{n}$ and define the following process.
+Choose a uniformly random point $x\in A$ and then moving to one of its $n$ neighbors with probability ${\rm P}_{p}(x,x\oplus e_{i})=\omega_{p}(\{x,x\oplus e_{i}\})/d_{p}(x)$, where $d_{p}(x)=\sum_{i=1}^{n}\omega_{p}(\{x,x\oplus e_{i}\})$ is the weighted degree of vertex $x$.
+Let $Y$ be the time that it takes the walk to exit $A$ for the first time.
+We refer to $\mathbb{E}_{p}[Y]$ as the mean first exit time of $A$ under the $p$-biased random walk.
+Corollary 1.5.
+Increasing subcubes maximize mean first exit time among all increasing subsets of the weighted hypercube $Q_{p}^{n}$ of the same cardinality.
+More precisely, for any increasing subset $A\subset\{0,1\}^{n}$,
+$\mathbb{E}_{p}[Y]\leq\frac{n}{\log_{p}(\mu_{p}(A))},$ (8)
+equality holds if and only if $g$ is indicator functions of increasing subcubes.
+1.2 Poincaré inequality on increasing sets
 Discrete-time random walks on finite graphs $G=(V,E)$ model stochastic processes where a walker iteratively transitions between adjacent vertices.
@@
 where $\|\cdot\|_{{\rm TV}}$ denotes the total variation, and $\varepsilon\in(0,1)$ governs approximation accuracy.
-An important example is the $n$-dimensional discrete cube.
-This is a graph $Q_{n}=(\{0,1\}^{n},E)$ with $2^{n}$ vertices indexed by $n$-dimensional Boolean vector.
-Two 0-1 vectors are adjacent if they differ in exactly one coordinate.
-We say a set $A\subseteq\{0,1\}^{n}$ is monotone if $x\in A$ and $x\leq y$ imply that $y\in A$.
-Here, $x\leq y$ denotes $x_{i}\leq y_{i}$ for all $i\in[n]$.
 Ding and Mossel [10] introduced the study of the random walk on $\{0,1\}^{n}$ censored to $A$.
 This walk, when at state $x\in A$, attempts to flip a uniformly random coordinate; it moves to the neighbor if it resides in $A$, otherwise it stays at $x$.
-Note that if we do not add some properties on $A$, the censored random walk may mix well or extremely poorly even when $A$ is large and connected, see Example 1 and 2 in [2].
-A significant open question (Question 1.1 in [1]) asked if the optimal mixing time $O_{c}(n\log n)$ holds for $\mu(A)\geq c$, where we use $\mu:=\mu_{1/2}^{\otimes n}$ to denote the uniform probability measure on the cube.
-Question 1.1 (Question 1.1 in [1]).
-Suppose that there exists a constant $c>0$ such that a monotone subset $A\subseteq\{0,1\}^{n}$ has measure $\mu(A)\geq c$.
+Note that if we do not impose any restrictions on $A$, the censored random walk may mix well, and also may mix extremely poorly even when $A$ is large and connected, see [17, Example 1,2].
+A significant open question ([10, Question 1.1]) asked if the optimal mixing time belongs to $O_{c}(n\log n)$ holds for increasing subsets $A$ with constant density under the uniform measure.
+Question 1.6 ([10, Question 1.1]).
+Suppose that there exists a constant $c>0$ such that a increasing subset $A\subseteq\{0,1\}^{n}$ has measure $\mu(A)\geq c$.
 Consider the censored random walk on $A$ defined above.
@@
 When the density $\mu(A)$ is constant, Ding and Mossel established a mixing time bound of $O(n^{3})$.
-Very recently, Fei and Pinto Jr.
-[2] show an improved mixing time bound of $O(n^{2})$ for constant-density monotone sets via establishing the following Poincaré inequality on the non-product space.
-Theorem 1.2 [2] Theorem 1.6).
-Let $A\subseteq\{0,1\}^{n}$ be a non-empty monotone set. We have for all $f:A\to\mathbb{R}$,
+Very recently, Fei and Ferreira Pinto Jr [17] show an improved mixing time bound of $O(n^{2})$ for constant-density increasing sets via establishing the following Poincaré inequality on the non-product space.
+Theorem 1.7 ([17, Theorem 1.6]).
+Let $A\subseteq\{0,1\}^{n}$ be a non-empty increasing set. It holds that for all $f:A\to\mathbb{R}$,
 ${\rm Var}_{A}[f]\leq\frac{1}{1-\sqrt{1-\mu(A)}}\cdot\mathcal{E}_{A}(f).$ (10)
-Here ${\rm Var}_{A}[f]$ stands for the variance of $f(x)$ where $x$ is a uniformly random element of $A$ and $\mathcal{E}_{A}(f)$ is called “restricted” Dirichlet form of a function on $A$ defined as follows
+Here ${\rm Var}_{A}[f]$ stands for the variance of $f(x)$ where $x$ is a uniformly random element of $A$, and $\mathcal{E}_{A}(f)$ is called “restricted” Dirichlet form of a function on $A$ defined as follows
 $\mathcal{E}_{A}(f)=\mathcal{E}_{A}(f,f):=\frac{1}{4}\cdot\sum\limits_{i=1}^{n}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right],$ (11)
-where the expectation is taken with respect to the uniform probability measure on the monotone subset $A\subseteq\{0,1\}^{n}$ and $x^{\oplus i}$ is obtained from $x$ by flipping the $i$th coordinate.
-Note that when $A=Q_{n}$, the above result recovers the classical Poincaré inequality on the hypercube.
+where the expectation is taken with respect to the uniform probability measure on the increasing subset $A\subseteq\{0,1\}^{n}$ and $x^{\oplus i}$ is obtained from $x$ by flipping the $i$th coordinate.
+Note that when $A=Q^{n}$, the above result recovers the classical Poincaré inequality on the hypercube.
 Their proof relies heavily on the connections between the directed heat flow and the directed Poincaré inequality.
-In this note, we provide a simple inductive proof of their Poincaré inequality, though yielding a slightly worse constant (theirs being 1 versus ours at 2).
-Nevertheless, we still obtain an $O(n^{2})$ mixing time bound for monotone sets $A$ of constant density via standard Markov chain theory (Theorem 12.4 in [3]).
-Specifically, we will prove the following theorem, a weaker version of Theorem 1.2.
+In this paper, we provide a simple inductive proof of their Poincaré inequality, though yielding a slightly worse constant (theirs being 1 versus ours at 2).
+Specifically, we prove the following theorem, a weaker version of Theorem 1.7.
 Theorem 1.8.
-Let $A\subseteq\{0,1\}^{n}$ be a non-empty monotone set. We have for all $f:A\to\mathbb{R}$,
+Let $A\subseteq\{0,1\}^{n}$ be a non-empty increasing set. We have for all $f:A\to\mathbb{R}$,
 ${\rm Var}_{A}[f]\leq\frac{2}{1-\sqrt{1-\mu(A)}}\cdot\mathcal{E}_{A}(f).$ (12)
-2 Preliminaries and Proofs
-In this section, we outline some notation and terminology, and present some simple lemmas for inductive proof of Theorem 1.3.
+Nevertheless, we still obtain an $O(n^{2})$ mixing time bound for increasing sets $A$ of constant density via standard Markov chain theory ([27, Theorem 12.4]).
+Corollary 1.9.
+Let $A\subseteq\{0,1\}^{n}$ be a non-empty increasing set. Then the random walk on $\{0,1\}^{n}$ censored to $A$ has mixing time
+$t_{{\rm mix}}\leq\frac{2n}{\mu(A)}\cdot\log(4\cdot 2^{n}\mu(A)).$
+Organization.
+This paper is organized as follows.
+Section 2 is devoted to our first main result: a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions.
+We develop the induction-on-dimension framework and verify the required two-point inequality in this setting.
+Section 3 turns to our second main result: an inductive proof of a Poincaré inequality on increasing subsets of the cube, which yields an $O(n^{2})$ upper bound on the mixing time of the censored random walk.
+The arguments once again follow the induction + five-point paradigm, though the analytic verification is much simple than the biased case.
+Finally, Section 4 presents concluding remarks and open problems.
+We highlight two directions for future research: (i) developing analytic proofs of weaker forms of Harper’s edge-isoperimetric inequality, and (ii) establishing suitable (possibly modified) log-Sobolev inequalities on increasing sets, which are closely related to the Ding–Mossel conjecture.
+2 Proof of Theorem 1.4
+2.1 Preliminaries
+In this subsection, we outline some notation and terminology, and present some simple lemmas for inductive proof of Theorem 1.4.
+For a parameter $p\in(0,1)$, the $p$-biased measure on $\{0,1\}$ assigns probability $p$ to $1$ and probability 1p1-p to 0.
+Let $\mu_{p}^{\otimes n}$ be the product distribution of $\mu_{p}$ over $\{0,1\}^{n}$.
+For $x\in\{0,1\}^{n}$, let $|x|=|\{i\in[n]:x_{i}=1\}|$ and then
+$\mu_{p}(x):=\mu_{p}^{\otimes n}(x)=p^{|x|}(1-p)^{n-|x|}.$
+We consider the space of real valued functions on $\{0,1\}^{n}$, equipped with the inner product
+$\left\langle f,g\right\rangle_{p}=\underset{x\sim\mu_{p}}{\mathbb{E}}[f(x)g(x)]=\int_{\{0,1\}^{n}}f(x)g(x)\mu_{p}(x)=\sum\limits_{x\in\{0,1\}^{n}}f(x)\cdot g(x)\mu_{p}(x).$
+Definition 2.1.
+For $q\geq 1$, the $q$-norm w.r.t. $p$-biased measure of $f:\{0,1\}^{n}\to\mathbb{R}$ is
+$\|f\|_{q,\mu_{p}}:=\left(\underset{x\sim\mu_{p}}{\mathbb{E}}[|f(x)|^{q}]\right)^{1/q}.$
+In the language of Boolean functions, there are two different ways to prove inequalities about $f:\{0,1\}^{n}\to\mathbb{R}$ by induction on $n$.
+One method, which might be called “induction by derivatives”, decomposes a function $f$ as follows: $f(x)=(-1)^{x_{i}}g_{1}(y)+g_{0}(y)$ for some functions $g_{0},g_{1}$ on the (n1)(n-1)-dimensional cube, where $y=(x_{1},\dots,x_{n-1})$.
+Such a decomposition is motivated from Fourier expansion.
+This induction method underlies classic proofs of Bonami’s lemma and the hypercontractive inequality on the hypercube [29].
+The other method, which we term “induction by restrictions”, proceeds via the restricted functions $g_{a}$ obtained by fixing the last coordinate to $a\in\{0,1\}$.
+One then expresses Dirichlet forms (or total influence) and variances recursively in terms of $g_{0}$ and $g_{1}$, applies the inductive hypothesis in dimension n1n-1, and controls the remaining cross-terms by verifying an explicit two-point (and occasionally four/five-point) inequality.
+This strategy—used in our proofs of Theorems 1.4 and 1.8—systematically reduces high-dimensional functional inequalities to finite-dimensional checks and often yields sharp constants.
+Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V$. Let
+$V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$
+and
+$A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}.$
+Let $\mu_{p}^{n-1}(A_{0})=a_{0},\mu_{p}^{n-1}(A_{1})=a_{1}$, and $\mu_{p}^{n}(A)=a=pa_{1}+(1-p)a_{0}$. Define $g_{0},g_{1}:(\{0,1\}^{n-1},\mu_{p}^{n-1})\to\mathbb{R}$ by
+$g_{0}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},0)\quad\text{and}\quad g_{1}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},1),$
+and the support of $g_{1}$ and $g_{0}$ are the whole set $A_{1}$ and $A_{0}$.
+Note that if $g$ is an increasing function, then the set $A$ is also an increasing set, since it is the support of an increasing function $g$.
+This implies $A_{0}\subset A_{1}$.
+Fact 2.2.
+If $A$ is an increasing set, then $A_{0}\subset A_{1}$.
+Proof.
+For any $x^{\prime}\in A_{0}$, we shall see that $x^{\prime}\in A_{1}$.
+Indeed, $\mathbbm{1}_{A_{0}}(x^{\prime})=1$ implies $(x^{\prime},0)\in A$ i.e.
+$\mathbbm{1}_{A}(x^{\prime},0)=1$.
+As $(x^{\prime},0)\leq(x^{\prime},1)$ and $A$ is increasing, $1=\mathbbm{1}_{A}(x^{\prime},0)\leq\mathbbm{1}_{A}(x^{\prime},1)=1$.
+∎
+By applying conditional expectation expansion to $n$th variable of Dirichlet forms, we can readily reduce the Dirichlet form of the original function to that of the restricted functions.
+Lemma 2.3.
+For any $g:\{0,1\}^{n}\to\mathbb{R}$, we have
+$\mathcal{E}_{p}^{n}(g,g)=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}.$ (13)
+Proof.
+Note that
+$\begin{split}\mathcal{E}_{p}^{n}(g,g)&=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\underset{x_{n}\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},x_{n})}(g(x^{\prime},x_{n})-g(y))^{2}\right]\\ &=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[p\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}+(1-p)\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}\right]\\ &=p\cdot\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in V_{1}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}+\sum\limits_{y\in V_{0}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}\right]\\ &+(1-p)\cdot\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in V_{1}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}+\sum\limits_{y\in V_{0}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}\right]\\ &=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+p\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\\ &+(1-p)\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}.\end{split}$
+We also need to use some simple properties of the function $f(t):=\frac{\log_{p}t}{t}$.
+Lemma 2.4.
+The function $f(t)$ is decreasing and convex for $t\in[0,1],p\in(0,1)$. Moreover, it satisfies the identity
+$f(\beta\cdot t)=\frac{\log_{p}(\beta\cdot t)}{\beta\cdot t}=\frac{1}{\beta}\cdot\frac{\log_{p}t}{t}+\frac{\log_{p}\beta}{\beta}\cdot\frac{1}{t}.$
+Proof.
+Note that $f^{\prime}(t)=\frac{1-\log t}{\log p\cdot t^{2}}\leq 0$ and $f^{\prime\prime}(t)=\frac{-3+2\log t}{\log p\cdot t^{3}}\geq 0$ when $t\in[0,1],p\in(0,1)$. ∎
+2.2 Induction Part
+We shall make several simple assumptions:
+  • We may assume without loss of generality that $g\geq 0$, since replacing $g$ by its absolute value preserves the right-hand side of (5), while it only decreases the left-hand side:
+    $\begin{split}\mathcal{E}_{p}(g,g)&=\underset{x\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}|g(x)-g(y)|^{2}\right]\\ &\geq\underset{x\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}(|g(x)|-|g(y)|)^{2}\right]=\mathcal{E}_{p}(|g|,|g|).\end{split}$
+  • We may assume the support of $g$ is the whole set $A$.
+Indeed, if we denote $B:={\rm supp}(g)=\{x\in\{0,1\}^{n}:g(x)>0\}$ with $B\subset A$.
+Then $\mathbb{E}_{p}[g]^{2}=\left(\sum_{x\in B}g(x)\mu_{p}(x)\right)^{2}$.
+Hence, replacing the support of $g$ with the whole set $A$ only increases the right-hand side of (5).
+It is easy to check that the theorem holds for $n=1$.
+Let $n\geq 2$, and suppose the statement of the theorem holds when $n$ is replaced by n1n-1.
+Using Lemma 2.3 to split the Dirichlet form of the original function to two restricted functions:
+$\mathcal{E}_{p}^{n}(g,g)=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}.$ (14)
+Here, we use the superscript $k$ with $k=n$ or n1n-1 to indicate that the Dirichlet form is the one defined on the $k$-dimensional hypercube.
+In the following, we adopt the same convention for other quantities, e.g., the measure $\mu_{p}$.
+By inductive assumption, we have
+$\begin{split}&p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})\geq\frac{\mathbb{E}_{p}[g_{1}]^{2}}{a_{1}}\log_{p}a_{1}\\ &p\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\geq\frac{\mathbb{E}_{p}[g_{0}]^{2}}{a_{0}}\log_{p}a_{0}.\end{split}$ (15)
+Thus our goal is to show that
+$\begin{split}p\mathcal{E}_{p}^{n}(g,g)&=p^{2}\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+p(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+p\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}\\ &\geq p\cdot\frac{\mathbb{E}_{p}[g_{1}]^{2}}{a_{1}}\log_{p}a_{1}+(1-p)\cdot\frac{\mathbb{E}_{p}[g_{0}]^{2}}{a_{0}}\log_{p}a_{0}+p\cdot\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}\\ &\underset{?}{\geq}\frac{\log_{p}(pa_{1}+(1-p)a_{0})}{pa_{1}+(1-p)a_{0}}\cdot\mathbb{E}_{p}[g]^{2}.\end{split}$ (16)
+  • Finally, we may assume both $g_{0},g_{1}$ to be constant on $A_{0}$ and on $A_{1}\setminus A_{0}$ (and of course $g_{i}$ vanish on $A_{i}^{c}$, in particular $g_{0}$ is 0 on $A_{1}\setminus A_{0}$).
+Indeed, if we replace $g_{j}$ with $\mathbb{E}_{p}[g_{j}]$ for $j=0,1$, then
+    $\begin{split}&\left\langle\mathbbm{1}_{A_{j}},\mathbb{E}_{x\sim\mu_{p}}[g_{j}]\right\rangle_{p}=\underset{x^{\prime}\sim\mu_{p}}{\mathbb{E}}[g_{j}(x^{\prime})\cdot\mathbbm{1}_{A_{j}}]\cdot a_{j}\leq\underset{x^{\prime}\sim\mu_{p}}{\mathbb{E}}[g_{j}(x^{\prime})\cdot\mathbbm{1}_{A_{j}}],\\ &\|\mathbb{E}_{x\sim\mu_{p}}[g_{1}]-\mathbb{E}_{x\sim\mu_{p}}[g_{0}]\|_{2,\mu_{p}}^{2}\leq\underset{x\sim\mu_{p}}{\mathbb{E}}[\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}]=\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2},\end{split}$
+    where the first inequality is due to $a_{j}=\mu_{p}^{n-1}(A_{j})\leq 1$ and the second inequality is due to Jensen’s inequality.
+Hence replacing $g_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of (16).
+Now let $\alpha$ be the value of $g_{1}$ on $A_{0}$ and $\gamma$ be the value of $g_{1}$ on $A_{1}\setminus A_{0}$. Let $\beta$ be the value of $g_{0}$ on $A_{0}$. Note that
+$\begin{split}&s_{1}:=\mathbb{E}_{p}[g_{1}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)=\alpha a_{0}+\gamma(a_{1}-a_{0}),\\ &s_{0}:=\mathbb{E}_{p}[g_{0}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)=\beta a_{0},\\ &\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &\quad\ =\sum\limits_{x^{\prime}\in A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})+\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})\\ &\quad\ =(\alpha-\beta)^{2}\cdot a_{0}+(a_{1}-a_{0})\cdot\gamma^{2}.\end{split}$
+Then
+[…]
+With the new notation, the inequality to be verified for the induction step is:
+$\begin{split}p\cdot\frac{\log_{p}a_{1}}{a_{1}}\cdot s_{1}^{2}&+(1-p)\cdot\frac{\log_{p}a_{0}}{a_{0}}\cdot s_{0}^{2}+p\cdot(\alpha-\beta)^{2}\cdot a_{0}+p\cdot(a_{1}-a_{0})\cdot\gamma^{2}\\ &\geq\frac{\log_{p}(pa_{1}+(1-p)a_{0})}{pa_{1}+(1-p)a_{0}}\cdot(ps_{1}+(1-p)s_{0})^{2}.\end{split}$ (17)
+Let $f(t):=\frac{\log_{p}t}{t}$ with $t\in[0,1]$. Then our goal is to prove that
+$ps_{1}^{2}f(a_{1})+(1-p)s_{0}^{2}f(a_{0})+pa_{0}(\alpha-\beta)^{2}+p(a_{1}-a_{0})\gamma^{2}\geq(ps_{1}+(1-p)s_{0})^{2}f(pa_{1}+(1-p)a_{0})).$ (18)
+Expressing $\beta$ and $\gamma$ as functions of $s_{i},a_{i}$ and of $\alpha$, LHS of the above is a quadratic in $\alpha$ with coefficients depending on $s_{i}$ and $a_{i}$: LHS $A\alpha^{2}+B\alpha+C$, where
+$\begin{split}&A:=p\left(a_{0}+\frac{a_{0}^{2}}{a_{1}-a_{0}}\right)\\ &B:=-2p\left(s_{0}+\frac{a_{0}s_{1}}{a_{1}-a_{0}}\right)\\ &C:=ps_{1}^{2}f(a_{1})+(1-p)s_{0}^{2}f(a_{0})+p\frac{s_{0}^{2}}{a_{0}}+p\frac{s_{1}^{2}}{a_{1}-a_{0}}.\end{split}$
+Minimizing LHS in $\alpha$ (when $\alpha=-\frac{B}{2A}=\frac{s_{0}a_{1}-s_{0}a_{0}+a_{0}s_{1}}{a_{1}a_{0}}$), we arrive, after some simple calculations, to the following inequality we need to verify:
+$ps_{1}^{2}f(a_{1})+(1-p)s_{0}^{2}f(a_{0})+p\frac{(s_{0}-s_{1})^{2}}{a_{1}}\geq(ps_{1}+(1-p)s_{0})^{2}f(pa_{1}+(1-p)a_{0})).$
+Next, let $T=\frac{s_{0}}{s_{1}}$. The above inequality transforms to a quadratic inequality in $T$:
+$pf(a_{1})+(1-p)f(a_{0})T^{2}+p\frac{(1-T)^{2}}{a_{1}}\geq(p+(1-p)T)^{2}f(pa_{1}+(1-p)a_{0})).$ (19)
+We need to check $h(T):=A^{\prime}T^{2}+B^{\prime}T+C^{\prime}\geq 0$ with the coefficients
+$\begin{split}&A^{\prime}:=(1-p)f(a_{0})+\frac{p}{a_{1}}-(1-p)^{2}f(pa_{1}+(1-p)a_{0}))\\ &B^{\prime}:=-\frac{2p}{a_{1}}-2p(1-p)f(pa_{1}+(1-p)a_{0}))\\ &C^{\prime}:=pf(a_{1})+\frac{p}{a_{1}}-p^{2}f(pa_{1}+(1-p)a_{0})).\end{split}$
+We shall verify $A^{\prime}\geq 0$ and $D=B^{\prime 2}-4A^{\prime}C^{\prime}\leq 0$, which will conclude the proof.
+Lemma 2.5.
+$A^{\prime}:=(1-p)f(a_{0})+\frac{p}{a_{1}}-(1-p)^{2}f(pa_{1}+(1-p)a_{0}))\geq 0$.
+Proof.
+Note that
+$\begin{split}A^{\prime}&=(1-p)f(a_{0})+\frac{p}{a_{1}}-(1-p)^{2}f(pa_{1}+(1-p)a_{0}))\\ &=\frac{(1-p)a_{1}f(a_{0})+p-(1-p)^{2}a_{1}f(pa_{1}+(1-p)a_{0}))}{a_{1}}.\end{split}$
+By Lemma 2.4, $f(pa_{1}+(1-p)a_{0}))\leq f(a_{0})$ and hence,
+$(1-p)^{2}a_{1}f(pa_{1}+(1-p)a_{0}))\geq-(1-p)^{2}a_{1}f(a_{0})\geq-(1-p)a_{1}f(a_{0}).$
+Then
+$A^{\prime}\geq\frac{(1-p)a_{1}f(a_{0})+p-(1-p)a_{1}f(a_{0})}{a_{1}}=\frac{p}{a_{1}}\geq 0.$
+It remains to verify the inequality $4A^{\prime}C^{\prime}\geq B^{\prime 2}$, which reduces to a two-point inequality:
+$pf(a_{1})+(1-p)f(a_{0})+(1-p)a_{1}f(a_{0})f(a_{1})\geq(p(1-p)a_{1}f(a_{0})+(1-p)^{2}a_{1}f(a_{1})+1)f(pa_{1}+(1-p)a_{0}).$ (20)
+2.3 Two-point Inequality Part
+Renaming the variables $x=a_{0}$ and $y=a_{1}$, and recalling the constraints on $a_{0}$ and $a_{1}$, we need to prove (20) for $0<x<y\leq 1$.
+$pf(y)+(1-p)f(x)+(1-p)yf(x)f(y)\geq(p(1-p)yf(x)+(1-p)^{2}yf(y)+1)f(py+(1-p)x)$ (21)
+Rearranging, this is easily seen to be equivalent to
+$\Delta_{p}(x,y)\geq\frac{p(1-p)^{2}y(f(x)-f(y))^{2}}{p(1-p)yf(x)+(1-p)^{2}yf(y)+1}$ (22)
+where
+$\Delta_{p}(x,y):=pf(y)+(1-p)f(x)-f(py+(1-p)x)$
+Note that $\Delta_{p}(x,y)\geq 0$ since $f(t):=\frac{\log_{p}t}{t}$ is convex.
+We now substitute $x=ty$ in $\Delta_{p}(x,y)$, with $0<t<1$, and expand it using Lemma 2.4. We have
+$\begin{split}\Delta_{p}(x,y)=\Delta_{p}(ty,y)&=pf(y)+(1-p)f(t\cdot y)-f((p+(1-p)t)\cdot y)\\ &=pf(y)+\frac{1}{t}\cdot(1-p)f(y)+\frac{1}{y}\cdot(1-p)f(t)-\frac{f(y)}{p+(1-p)t}-\frac{f(p+(1-p)t)}{y}\\ &=\frac{p(1-p)(t-1)^{2}}{t(p+(1-p)t)}\cdot f(y)+\frac{1}{y}\cdot\left(\frac{(1-p)\log_{p}t}{t}-\frac{\log_{p}(p+(1-p)t)}{p+(1-p)t}\right)\end{split}$
+As to the right-hand side of (22), we have
+$\begin{split}{\rm RHS}(x,y)&={\rm RHS}(ty,y)=\frac{p(1-p)^{2}y[(\frac{1}{t}-1)f(y)+\frac{1}{y}\cdot\frac{\log_{p}t}{t}]^{2}}{[\frac{p(1-p)}{t}+(1-p)^{2}]yf(y)+p(1-p)\cdot\frac{\log_{p}t}{t}+1}\\ &=\frac{y[(\frac{1}{t}-1)f(y)+\frac{1}{y}\cdot\frac{\log_{p}t}{t}]^{2}}{[\frac{1}{t(1-p)}+\frac{1}{p}]yf(y)+\frac{\log_{p}t}{(1-p)t}+\frac{1}{p(1-p)^{2}}}.\end{split}$
+Taking $z:=yf(y)$,
+$\Delta_{p}(x,y)\geq{\rm RHS}\Leftrightarrow y\cdot\Delta_{p}(x,y)\geq y\cdot{\rm RHS}\Leftrightarrow Az+B\geq\frac{(Cz+D)^{2}}{Ez+F},$
+where
+$z=yf(y)=\log_{p}(y)>0\ \text{as $0<y<1$}$
+and $A,B,\dots,F$ depend only on $t$ and $p$. Specifically,
+$\begin{cases}A=\frac{p(1-p)(t-1)^{2}}{t(p+(1-p)t)},\\ B=\frac{(1-p)\log_{p}t}{t}-\frac{\log_{p}(p+(1-p)t)}{p+(1-p)t},\\ C=\frac{1-t}{t},\\ D=\frac{\log_{p}t}{t},\\ E=\frac{1}{t(1-p)}+\frac{1}{p}\geq 0,\\ F=\frac{\log_{p}t}{(1-p)t}+\frac{1}{p(1-p)^{2}}\geq 0.\end{cases}$
+Thus we need to show that
+$(Az+B)(Ez+F)\geq(Cz+D)^{2}.$
+Observe that
+$AE=\frac{(t-1)^{2}}{t^{2}}=C^{2}.$
+Therefore, this reduces to a linear inequality in $z$:
+$(AF+BE-2CD)\cdot z\geq D^{2}-BF.$
+This holds for all non-negative $z$ if and only if $AF+BE\geq 2CD$ and $BF\geq D^{2}$.
+Lemma 2.6.
+For any $t\in(0,1)$ and $p\in(0,1)$ we have $AF+BE\geq 2CD$.
+Proof.
+Simplifying and rearranging terms, the inequality reduces to verifying that
+$F(t,p):=(1-p)t\log_{p}t+p(t-1)^{2}-(p+(1-p)t)\log_{p}(p+(1-p)t)\geq 0.$
+We compute the derivatives of $F(t,p)$ with respect to $t$:
+$\begin{split}&\frac{\partial}{\partial t}F(t,p)=\frac{1-p}{\ln(p)}\cdot\ln\left(\frac{t}{p+(1-p)t}\right)+2p\cdot(t-1),\\ &\frac{\partial^{2}}{\partial t^{2}}F(t,p)=\frac{p(1-p)}{\ln(p)}\cdot\frac{1}{t(p+(1-p)t)}+2p,\\ &\frac{\partial^{3}}{\partial t^{3}}F(t,p)=\frac{p(1-p)}{\ln(1/p)}\cdot\frac{p+2(1-p)t}{t^{2}(p+(1-p)t)^{2}}\geq 0.\end{split}$
+Hence $\frac{\partial}{\partial t}F(t,p)$ is a convex function of $t$, implying that it can vanish at most twice on the interval $t\in(0,1)$.
+Note that $\frac{\partial}{\partial t}F(t,p)|_{t\to 0}>0$, $\frac{\partial}{\partial t}F(t,p)|_{t=1}=0$.
+Therefore, there exists $t_{0}:=t_{0}(p)\in(0,1)$ such that $\frac{\partial}{\partial t}F(t,p)|_{t=t_{0}}=0$ and $\frac{\partial}{\partial t}F(t,p)\geq 0$ for $t\in(0,t_{0}]$ and $F(t,p)\leq 0$ for $t\in(t_{0},1)$.
+Since $F(0,p)=F(1,p)=0$, the function $F(t,p)$ first increases from 0 at $t=0$ and then decreases back to 0 at $t=1$.
+Consequently, $F(t,p)\geq 0$ for all $t\in(0,1)$, which establishes the desired inequality.
+∎
+Lemma 2.7.
+For any $t\in(0,1)$ and $p\in(0,1)$ we have $BF\geq D^{2}$.
+Proof.
+Simplifying and rearranging terms, this inequality reduces to show that
+$G(t,p):=(1-p)(p+(1-p)t)\log_{p}t-p(1-p)\log_{p}(t)\cdot\log_{p}(p+(1-p)t)-t\cdot\log_{p}(p+(1-p)t)\geq 0.$
+Clearly, $G(0,p)=G(1,p)=0$.
+We aim to show that $G(t,p)$ increases from 0 at $t=0$, attains its maximum at some interior point, and then decreases back to 0 at $t=1$; hence $G(t,p)\geq 0$ for all $t\in(0,1)$.
+We begin by computing the derivative:
+$\begin{split}\frac{\partial}{\partial t}G(t,p)&=\frac{1}{\ln(p)}\left(\frac{p(1-p)}{t}+(1-p)^{2}\ln(t)+(1-p)^{2}-\frac{p(1-p)}{\ln(p)}\cdot\frac{\ln(p+(1-p)t)}{t}\right)\\ &-\frac{1}{\ln(p)}\left(\frac{p(1-p)^{2}}{\ln(p)}\cdot\frac{\ln(t)}{p+(1-p)t}+\ln(p+(1-p)t)+\frac{(1-p)t}{p+(1-p)t}\right).\end{split}$
+Define $H(t,p):=t(p+(1-p)t)\cdot\frac{\partial}{\partial t}G(t,p)$. To analyze its sign, we consider successive derivatives of $H$.
+$\begin{split}&\frac{\partial}{\partial t}H(t,p)=\frac{1}{\ln(p)^{2}}\left[-(p-1)^{2}p\left(2+\ln(t)+\ln(p+t-pt)\right)\right]\\ &+\frac{1}{\ln(p)}\left[(-2t+p(2t-1)\ln(p+t-pt)\right]\\ &-\frac{1}{\ln(p)}\left[3(p-1)p(1+p(t-1)-2t)+(p-1)^{2}(-2t+p(2t-1))\ln(t)\right];\\ &\frac{\partial^{2}}{\partial t^{2}}H(t,p)=\frac{1}{\ln(p)^{2}}\left(-(p-1)^{2}p\left(\frac{1}{t}+\frac{1-p}{p+t-pt}\right)\right)\\ &+\frac{1}{\ln(p)}\left(-3(-2+p)(p-1)p-\frac{(p-1)^{2}(-2t+p(2t-1))}{t}+\frac{(1-p)(-2t+p(2t-1))}{p+t-pt}\right)\\ &+\frac{1}{\ln(p)}\left(-2(p-1)^{3}\ln(t)+2(p-1)\ln(p+t-pt)\right);\\ &\frac{\partial^{3}}{\partial t^{3}}H(t,p)=-\frac{(p-1)^{2}p}{t^{2}(p+t-pt)^{2}\ln(p)^{2}}\cdot h(t,p),\end{split}$
+where $h(t,p):=-2t^{2}+2pt(-1+2t)+p^{2}(-1+2t-2t^{2})+(-6p(-1+t)t^{2}+4t^{3}+p^{2}(t-1)^{2}(1+2t))\ln(p)$.
+Thus it suffices to consider the sign of function $h(t,p)$ for $t\in(0,1)$ and $p\in(0,1)$.
+Note that
+$\begin{split}&\frac{\partial}{\partial t}h(t,p)=-2(p-1)(-2t+p(-1+2t))+6(p-2)(p(t-1)-t)t\ln(p),\\ &\frac{\partial^{2}}{\partial t^{2}}h(t,p)=-4(p-1)^{2}+6(p-2)(-2t+p(-1+2t))\ln(p),\\ &\frac{\partial^{3}}{\partial t^{3}}h(t,p)=12(p-2)(p-1)\ln(p)<0.\end{split}$
+Hence $\frac{\partial^{2}}{\partial t^{2}}h(t,p)$ is strictly decreasing on $(0,1)$. Evaluating at $t\to 0$, we obtain
+$\lim_{t\to 0}\frac{\partial^{2}}{\partial t^{2}}h(t,p)=-4(p-1)^{2}-6(p-2)p\ln(p).$
+Write $f(p):=-4(p-1)^{2}-6(p-2)p\ln(p)$. Note that f(0)=4,f(1)=0f(0)=-4,f(1)=0, and
+$\begin{split}&f^{\prime}(p)=-2(-10+7p+6(p-1)\ln(p)),\\ &f^{\prime\prime}(p)=-26+\frac{12}{p}-12\ln(p),\\ &f^{\prime\prime\prime}(p)=-\frac{12(1+p)}{p^{2}}<0.\end{split}$
+Therefore $f^{\prime\prime}(p)$ has a unique zero $p_{0}\in(0,1)$ since $f^{\prime\prime}(0)\to\infty$ and $f^{\prime\prime}(1)=-14$.
+It follows that $f^{\prime}(p)$ increases on $(0,p_{0})$ and decreases on $(p_{0},1)$.
+As $f^{\prime}(0)\to$-$\infty$ and $f^{\prime}(1)=6$, we deduce that $f(p)$ first decreases on $(0,p_{0})$ and then increases on $(p_{0},1)$, so $f(p)=\lim_{t\to 0}\frac{\partial^{2}}{\partial t^{2}}h(t,p)\leq 0$ for all $p\in(0,1)$.
+Consequently, $\frac{\partial^{2}}{\partial t^{2}}h(t,p)$ is negative, and since $\frac{\partial}{\partial t}h(t,p)\leq\frac{\partial}{\partial t}h(t,p)|_{t=0}=2p(p-1)<0$, we have $h(t,p)<0$ throughout $(0,1)$.
+Thus, $\frac{\partial^{3}}{\partial t^{3}}H(t,p)>0$.
+Next, note that $\lim_{t\to 0}\frac{\partial^{2}}{\partial t^{2}}H(t,p)\to$-$\infty$ and $\frac{\partial^{2}}{\partial t^{2}}H(t,p)|_{t=1}>0$ with $\frac{\partial^{2}}{\partial t^{2}}H(t,p)$ increasing on $t\in(0,1)$.
+Hence it has a unique zero $t_{0}\in(0,1)$, and $\frac{\partial}{\partial t}H(t,p)$ decreases on $(0,t_{0})$ then increases on $(t_{0},1)$.
+Since $\frac{\partial}{\partial t}H(t,p)|_{t=0},\frac{\partial}{\partial t}H(t,p)|_{t=1}>0$ and $\frac{\partial}{\partial t}H(t,p)|_{t=t_{0}}<0$, there exists $0<t_{1}<t_{2}<1$ such that $H(t,p)$ first increases on $(0,t_{1})$, then decreases on $t_{1},t_{2}]$ and finally increases again on $(t_{2},1)$.
+Because $H(0,p)=H(1,p)=0$, there exists $t^{\prime}\in[0,1]$ such that $H(t,p)>0$ for $t\in(0,t^{\prime})$ and $H(t,p)\leq 0$ for $t\in[t^{\prime},1)$.
+It implies that $G(t,p)$ increases on $t\in(0,t^{\prime})$ and then deceases in $t\in[t^{\prime},1)$.
+Since $G(0,p)=G(1,p)=0,$ we conclude that $G(t,p)\geq 0$ for all $t\in[0,1]$.
+3 Proof of Theorem 1.8
+3.1 Preliminaries
+In this subsection, we outline some notation and terminology, and present some simple lemmas for inductive proof of Theorem 1.8.
 Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V^{n}$. Let
@@
 Let $a_{0}:=\mu^{n-1}(A_{0}),a_{1}:=\mu^{n-1}(A_{1})$, and note that
-$\begin{split}a&:=\mu^{n}(A)=\frac{1}{2^{n}}\sum_{x\in V^{n}}\mathbbm{1}_{A}(x)=\frac{1}{2}\cdot\frac{1}{2^{n-1}}\left(\sum_{(x^{\prime},1)\in V_{1}}\mathbbm{1}_{A}(x^{\prime},1)+\sum_{(x^{\prime},0)\in V_{0}}\mathbbm{1}_{A}(x^{\prime},0)\right)\\ &=\frac{1}{2}\cdot\frac{1}{2^{n-1}}\left(\sum_{x^{\prime}\in V^{n-1}}\mathbbm{1}_{A_{1}}(x^{\prime})+\sum_{x^{\prime}\in V^{n-1}}\mathbbm{1}_{A_{0}}(x^{\prime})\right)=\frac{1}{2}\mu^{n-1}(A_{1})+\frac{1}{2}\mu^{n-1}(A_{0})=\frac{a_{0}+a_{1}}{2}.\end{split}$ (5)
+$\begin{split}a:=\mu^{n}(A)=\frac{1}{2}\mu^{n-1}(A_{1})+\frac{1}{2}\mu^{n-1}(A_{0})=\frac{a_{0}+a_{1}}{2}.\end{split}$ (23)
 Define $f_{0}:A_{0}\to\mathbb{R},\ f_{1}:A_{1}\to\mathbb{R}$ by
 $f_{0}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},0)\quad\text{and}\quad f_{1}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},1).$
-Fact 2.1.
-If $A$ is an increasing set, then $A_{0}\subset A_{1}$.
-Proof.
-For any $x^{\prime}\in A_{0}$, we shall see that $x^{\prime}\in A_{1}$. It is easy to see because $(x^{\prime},0)\leq(x^{\prime},1)$ and $A$ is increasing. ∎
-By applying conditional expectation expansion to $n$th variable of restricted Dirichlet forms, we can readily reduce the Dirichlet form of the original function to that of the restricted functions.
+By expanding the Dirichlet form with respect to the conditional expectation on the $n$th variable, we can readily express the Dirichlet form of the original function in terms of those of its restricted versions.
 Lemma 3.1.
@@
 Proof.
-We use $x^{\prime}$ to denote the first $(n-1)$-coordinates of $x$. Note that
-$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\sum\limits_{i=1}^{n-1}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right]+\frac{1}{4}\cdot\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus n}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1-x_{n})\in A\}}\right]\\ &=\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},1)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},0)\in A\}}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1)\in A\}}\right]\\ &=\frac{a_{1}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{1}\}}\right]+\frac{a_{0}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{0}\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right],\end{split}$
+We use $x^{\prime}$ to denote the first (n1)(n-1)-coordinates of $x$. Note that
+$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\sum\limits_{i=1}^{n-1}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right]+\frac{1}{4}\cdot\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus n}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1-x_{n})\in A\}}\right]\\ &=\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},1)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1)\in A\}}\right]\\ &=\frac{a_{1}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{1}\}}\right]\\ &+\frac{a_{0}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{0}\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right],\end{split}$
 where in the second equality, we observe that
 $\begin{split}\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}&\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\right]=\frac{1}{|A|}\cdot\sum\limits_{x^{\prime}\in A_{1}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\\ &=\frac{1}{|A|}\cdot\sum\limits_{x^{\prime}\in A_{0}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}=\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right].\end{split}$
-Similarly, we can reduce the variance of the original function to that of the restricted functions.
+Similarly, by conditioning on the $n$th coordinate, the variance of the original function can be expressed in terms of the variances of its restricted functions.
 Lemma 3.2.
@@
 Let $\mu=\underset{x\sim A}{\mathbb{E}}[f(x)],\mu_{1}=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]$ and $\mu_{0}=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]$. Then
-$\begin{split}\mu&=\frac{1}{|A|}\sum_{x\in A}f(x)=\frac{1}{|A|}\sum_{x^{\prime}\in A_{1}}f(x^{\prime},1)+\frac{1}{|A|}\sum_{x^{\prime}\in A_{0}}f(x^{\prime},0)=\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]+\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\\ &=\frac{a_{1}}{2a}\mu_{1}+\frac{a_{0}}{2a}\mu_{0}.\end{split}$ (8)
-Thus
-$\begin{split}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]&=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\frac{a_{1}}{2a}\mu_{1}-\frac{a_{0}}{2a}\mu_{0}\right|^{2}\right]=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}+\mu_{1}-\frac{a_{1}}{2a}\mu_{1}-\frac{a_{0}}{2a}\mu_{0}\right|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}+\frac{a_{0}}{2a}(\mu_{1}-\mu_{0})\right|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[\left|f_{1}-\mu_{1}\right|^{2}\right]+2\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}-\mu_{1})\cdot\frac{a_{0}}{2a}(\mu_{1}-\mu_{0})\right]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}\\ &={\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}.\end{split}$
-Similarly, we have
+$\mu=\frac{1}{|A|}\sum_{x\in A}f(x)=\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]+\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]=\frac{a_{1}}{2a}\mu_{1}+\frac{a_{0}}{2a}\mu_{0}.$ (26)
+Thus by the formula $\mathbb{E}[(f-b)^{2}]={\rm Var}[f]+(\mathbb{E}[f]-b)^{2}$, we have that
+$\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]={\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2},$
+and
 $\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]={\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}.$
 Thus we have
-$\begin{split}{\rm Var}_{A}[f]&=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{1}a_{0}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}^{2}+a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(\mu_{1}-\mu_{0})^{2}\end{split}$
-3 An inductive proof of Theorem 1.3
-Our goal is to prove that for all $f:A\to\mathbb{R}$,
+$\begin{split}{\rm Var}_{A}[f]&=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{1}a_{0}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(\mu_{1}-\mu_{0})^{2}\end{split}$
+3.2 Induction part
+Our goal is to prove that for every function $f:A\to\mathbb{R}$,
 $\mathcal{E}_{A}(f)\geq c\cdot(1-\sqrt{1-a})\cdot{\rm Var}_{A}[f],$ (27)
-where $c>0$ is a universal constant (we can take $c=0.5$).
-It is easy to check that the theorem holds for $n=1$.
-Let $n\geq 2$, and suppose the statement of the theorem holds when $n$ is replaced by $n-1$.
-Using Lemma 2.2 to split the Dirichlet form of the original function to two restricted functions:
+where $c>0$ is a universal constant (we may take $c=0.5$).
+It is easy to verify that the theorem holds for $n=1$.
+Let $n\geq 2$, and suppose the statement of the theorem holds when $n$ is replaced by n1n-1.
+Applying Lemma 3.1 to split the Dirichlet form of the original function into two restricted functions:
 $\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$ (28)
@@
 Thus our goal is to show that
-$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\geq\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}(c-c\sqrt{1-a}){\rm Var}_{A}[f].\end{split}$ (12)
+$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\geq\frac{a_{1}}{2a}(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}(c-c\sqrt{1-a}){\rm Var}_{A}[f].\end{split}$ (30)
 We also use Lemma 3.2 to split the variance of the original function to two restricted functions:
-${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}^{2}+a_{0}a_{1}^{2}}{8a^{3}}\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (13)
+${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (31)
 Thus it suffices to show that
-$\begin{split}&\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(c-c\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (14)
+$\begin{split}&\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}\frac{a_{1}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}(c-c\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (32)
 It is equivalent to show that
-$\begin{split}&ca_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+ca_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}c\frac{a_{1}a_{0}}{a_{0}+a_{1}}(1-\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (15)
+$\begin{split}&ca_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+ca_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]\\ &+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\underset{?}{\geq}c\frac{a_{1}a_{0}}{a_{0}+a_{1}}(1-\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (33)
 Lemma 3.3.
@@
 and
-$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}(x^{\prime})-\mu_{0}-(f_{1}(x^{\prime})-\mu_{1})+\mu_{0}-\mu_{1}|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}-\mu_{0}|^{2}\right]+\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{1}-\mu_{1}|^{2}\right]+2\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]+(\mu_{0}-\mu_{1})^{2}+2(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (17)
+$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}(x^{\prime})-\mu_{0}-(f_{1}(x^{\prime})-\mu_{1})+\mu_{0}-\mu_{1}|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}-\mu_{0}|^{2}\right]+\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{1}-\mu_{1}|^{2}\right]+2\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &+(\mu_{0}-\mu_{1})^{2}+2(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (35)
 For the sake of simplicity, we denote $C_{1}=\frac{1}{2}\cdot a_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}})$ and $C_{0}=\frac{1}{2}\cdot a_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}})$. Then
@@
 Now we focus on the term
-$\begin{split}&\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &=\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)|\mu_{1}-f_{1}(x^{\prime})|^{2}+\left(C_{0}+\frac{a_{0}}{2}\right)|f_{0}(x^{\prime})-\mu_{0}|^{2}+a_{0}(f_{0}(x^{\prime})-\mu_{0})(\mu_{1}-f_{1}(x^{\prime}))\\ &=\frac{1}{|A_{0}|}\sum_{x^{\prime}\in A_{0}}\vec{b}_{x^{\prime}}^{\top}G_{x^{\prime}}\vec{b}_{x^{\prime}},\end{split}$ (19)
+$\begin{split}&\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &=\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)|\mu_{1}-f_{1}(x^{\prime})|^{2}+\left(C_{0}+\frac{a_{0}}{2}\right)|f_{0}(x^{\prime})-\mu_{0}|^{2}\\ &+a_{0}(f_{0}(x^{\prime})-\mu_{0})(\mu_{1}-f_{1}(x^{\prime}))\\ &=\frac{1}{|A_{0}|}\sum_{x^{\prime}\in A_{0}}\vec{b}_{x^{\prime}}^{\top}G_{x^{\prime}}\vec{b}_{x^{\prime}},\end{split}$ (37)
 where
-$\vec{b}_{x^{\prime}}=\begin{bmatrix}f_{0}(x^{\prime})-\mu_{0}\\ \mu_{1}-f_{1}(x^{\prime})\end{bmatrix},\\ G_{x^{\prime}}=\begin{bmatrix}a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}&\frac{a_{0}}{2}\\ \frac{a_{0}}{2}&a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}.\end{bmatrix}\$
+$\vec{b}_{x^{\prime}}=\begin{bmatrix}f_{0}(x^{\prime})-\mu_{0}\\ \mu_{1}-f_{1}(x^{\prime})\end{bmatrix},\\ G_{x^{\prime}}=\begin{bmatrix}a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}&\frac{a_{0}}{2}\\ \frac{a_{0}}{2}&a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}\end{bmatrix}$.\\
 We claim that the matrix $G_{x^{\prime}}$ is positive semidefinite.
@@
 $\frac{a_{0}^{2}}{4}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-\frac{a_{0}^{2}}{4}\geq 0.$ (38)
-Let $s:=\sqrt{1-a_{0}}$ and $t:=\sqrt{1-a_{1}}$, then $0\leq t\leq s\leq 1$, $21-a=(s^{2}+t^{2})/$ 2 and
+Let $s:=\sqrt{1-a_{0}}$ and $t:=\sqrt{1-a_{1}}$, then $0\leq t\leq s\leq 1$, $1-a=(s^{2}+t^{2})/2$ and
 $\begin{split}&(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-1\\ &=\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-s+1\right)\cdot\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-t+1\right)-1\\ &=-\frac{1}{2}\cdot(s+t-2)\cdot(\sqrt{2}\sqrt{s^{2}+t^{2}}-s-t)\geq 0,\end{split}$ (39)
@@
 Hence replacing $f_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of (33). ∎
+3.3 Five-point Inequality Part
 Now let $\alpha$ be the value of $f_{1}$ on $A_{0}$ and $\beta$ be the value of $f_{1}$ on $A_{1}\setminus A_{0}$. Let $\gamma$ be the value of $f_{0}$ on $A_{0}$. Then we compute
@@
 ${\rm Var}_{A_{0}}[f_{0}]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}^{2}]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}]^{2}=\frac{a_{0}\cdot 2^{n-1}\cdot\gamma^{2}}{2^{n-1}\cdot a_{0}}-\left(\frac{a_{0}\cdot 2^{n-1}\cdot\gamma}{2^{n-1}\cdot a_{0}}\right)^{2}=0.$ (41)
-Thus our goal is to prove that
+Thus to verify (33), it suffices to prove the following five-point inequality with $c=\frac{1}{2}$.
 $\begin{split}&c(\sqrt{1-a}-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})\cdot(\alpha-\beta)^{2}+\frac{a_{1}}{2}\cdot(\gamma-\alpha)^{2}\\ &\geq\frac{c}{a_{1}+a_{0}}\cdot(1-\sqrt{1-a})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}.\end{split}$ (42)
@@
 Then it suffice to prove that
-$\begin{split}&\frac{c\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})}+\frac{a_{1}}{2}\cdot(\alpha-\gamma)^{2}\geq\frac{c}{2(1+\sqrt{1-a})}\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\\ &\Leftrightarrow c\cdot(1+\sqrt{1-a})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\geq 0.\end{split}$ (27)
+$\begin{split}&\frac{c\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})}+\frac{a_{1}}{2}\cdot(\alpha-\gamma)^{2}\geq\frac{c}{2(1+\sqrt{1-a})}\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\\ &\Longleftrightarrow c\cdot(1+\sqrt{1-a})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\geq 0.\end{split}$ (45)
 Note that
 $a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)=(a_{0}-a_{1})\cdot(\alpha-\beta)+a_{1}\cdot(\alpha-\gamma)$
-After some simple calculations, the following inequality we need to verify:
-$\begin{split}&c\cdot(1-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\cdot(a_{0}-a_{1})\cdot(\alpha-\beta)\cdot(\alpha-\gamma)\geq 0.\end{split}$ (28)
-If $\alpha=\gamma$, then it is trivially true.
-Thus we can assume that $\alpha\neq\gamma$ and let $T:=\frac{\alpha-\beta}{\alpha-\gamma}\cdot(a_{1}-a_{0})$.
-Then the above inequality transforms to a quadratic inequality in $T$:
+After some simple calculations, the inequality to be verified becomes
+$\begin{split}&c(1-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\cdot(a_{0}-a_{1})\cdot(\alpha-\beta)\cdot(\alpha-\gamma)\geq 0.\end{split}$ (46)
+If $\alpha=\gamma$, the inequality is trivially satisfied.
+Thus, assume $\alpha\neq\gamma$ and set $T:=\frac{\alpha-\beta}{\alpha-\gamma}\cdot(a_{1}-a_{0})$.
+Then the above inequality can be rewritten as a quadratic in $T$:
 $\begin{split}&c\cdot(1-\sqrt{1-a_{1}})\cdot T^{2}+2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot T\\ &a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\geq 0\end{split}$ (47)
-Now let us rewrite the Left-hand side of the above inequality as a quadratic in $T$ with coefficients depending only on $a_{1},a_{0}$:
+We may rewrite the left-hand side as a quadratic polynomial in $T$ ($AT^{2}+BT+C$), whose coefficients depend only on $a_{1},a_{0}$:
 $\begin{split}&A:=c\cdot(1-\sqrt{1-a_{1}})\geq 0\\ &B:=2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\\ &C:=a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}}).\end{split}$
-It only needs to verify $\Delta=B^{2}-4AC\leq 0$ with $c=0.5$, which will conclude the proof.
+To ensure the quadratic is nonnegative for all $T$, it suffices to check that its discriminant satisfies $\Delta=B^{2}-4AC\leq 0$ with $c=0.5$, which will conclude the proof.
 Lemma 3.4.
@@
 $\begin{split}&y-2+2\sqrt{1-y}\leq 0\\ &\Leftrightarrow 2\sqrt{1-y}\leq 2-y\\ &\Leftrightarrow y^{2}\geq 0.\end{split}$ (49)
+4 Conclusion and open problems
+The weak Harper’s edge-isoperimetric inequality (2) has several easy proofs based on induction.
+Once using conditional expectation to expand the total influence and by inductive assumption, it remains to check that the following two-point inequality for $a_{0},a_{1}\in(0,1)$
+$a_{1}\log_{2}\left(\frac{1}{a_{1}}\right)+a_{0}\log_{2}\left(\frac{1}{a_{0}}\right)+|a_{1}-a_{0}|\geq(a_{1}+a_{0})\log_{2}\left(\frac{2}{a_{1}+a_{0}}\right),$
+which is not hard to verify, using, e.g., concavity of the entropy function $H(x)=-x\log x-(1-x)\log(1-x)$ on $[0,1]$.
+It may be more amazing that we can prove the ‘sharp’ Harper’s edge-isoperimetric inequality by using entropy method, see Section 4.4 in [7].
+Since numerous classical results in Boolean function analysis–including the hypercontractivity inequality–can be established via entropy methods, see for example [32], we are very interested in whether other methods, especially analytical methods (or spectral graph theory), can be used to prove tight Harper’s edge-isoperimetric inequality.
+A natural next target beyond our $O(n^{2})$ mixing bound for constant-density increasing sets is the full Ding–Mossel conjecture asking for an $O(n\log n)$ mixing time for the censored walk when $\mu(A)\geq\varepsilon>0$ ([10, Question 1.1]).
+As observed in [17], a route to such an $O(n\log n)$ bound is to establish an appropriate (standard or modified) log-Sobolev inequality for the censored chain on $A$.
+Following the general relations between log-Sobolev and mixing time bound (e.g.,[6, 9]), it is natural to ask what the correct dependence on the density $\mu(A)$ should be in any valid log–Sobolev (or modified log-Sobolev) inequality for increasing sets.
+Our inductive two-point machinery is well suited for quadratic (L2) control: it converts high-dimensional Dirichlet forms and variances into lower-dimensional analogues plus a finite list of analytics inequalities (e.g., two-point inequalities) that one can verify exactly.
+However, in log-Sobolev inequalities, the relative entropy is a nonlinear, non-quadratic functional—this introduces challenges when applying the induction method to control it.
+Hence, to establish log-Sobolev inequalities in non-product spaces, further developing the induction method or looking for new methods is needed.

### v3

@@
-Functional inequalities on the biased and restricted cube: an induction approach
+Functional Inequalities and Random Walks on Increasing Subsets of the Hypercube
 Abstract
-We develop new discrete functional inequalities on the hypercube via the induction-by-restrictions method.
-This method reduces high-dimensional inequalities to explicit low-dimensional analytic verifications and has recently proved effective for many discrete functional inequalities.
-We establish two results along this framework.
-First, we prove a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions, which recovers the classic biased edge-isoperimetric inequality for increasing sets and identifies increasing subcubes as the extremizers.
+Motivated by random walks on subsets of the hypercube, we prove two discrete functional inequalities on the hypercube.
+First, we give a short, elementary proof of the Poincaré inequality on increasing subsets of the cube recently established by Fei and Ferreira Pinto Jr [fei2025spectral], which yields an $O(n^{2})$ upper bound on the mixing time of censored random walks, improving upon previous bounds.
+Second, adapting Samorodnitsky’s induction method [Samorodnitsky17] to the $p$-biased setting, we establish a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions, which recovers the classic biased edge-isoperimetric inequality for increasing sets and identifies increasing subcubes as the extremizers.
 This result also admits a probabilistic interpretation in terms of maximizing the mean first exit time of biased random walks.
-Second, we give an inductive proof of a Poincaré inequality on increasing subsets of the cube that was recently established by Fei and Ferreira Pinto Jr [17], yielding an $O(n^{2})$ upper bound on the mixing time of censored random walks, improving upon previous bounds.
+Keywords: Random walks, mixing time, Poincaré inequality, edge-isoperimetry, $p$-biased hypercube, mean first exit time
+MSC codes: 06E30, 05C81, 60C05, 60G50
 1 Introduction
-Functional inequalities on the discrete cube, such as the Poincaré, log-Sobolev, and edge-isoperimetric inequalities, form a central part of modern discrete analysis.
+Functional inequalities on the discrete hypercube/Hamming graph, such as the Poincaré, log-Sobolev, and edge-isoperimetric inequalities, form a central part of modern discrete analysis.
 They connect Boolean function analysis, discrete isoperimetric problems, and the spectral theory of Markov chains, providing powerful tools for studying concentration of measure, threshold phenomena, and mixing times of random walks.
-In the classical uniform product setting on $\{0,1\}^{n}$, these inequalities are well understood: sharp constants are known (e.g., the exact log-Sobolev constant [6]) and elegant proofs arise via tensorization [20], semigroup methods [23, 24], or discrete Fourier analysis [13].
-However, once one departs from the product setting–by restricting to structured subsets or by working under biased measures–classical approaches often fail or lose sharpness.
-Recent progress highlights both the challenges and the opportunities in these non-product environments: FKN-type structure theorems and sharp junta thresholds for low-degree Boolean functions on the slice/Johnson graph [18], (near) sharp log-Sobolev inequalities on the multislice [19], and representation-theoretic analyses of low-degree functions on the symmetric group that yield asymptotically sharp edge-isoperimetric bounds for the transposition graph [14].
-These developments motivate the search for new discrete functional inequalities adapted to non-product environments.
-A case in point that illustrates both the promise and the limitations of classical tools is the following.
-On the uniform cube, the logarithmic Sobolev inequality is tight for general real-valued functions, yet when specialized to indicator functions it recovers the edge-isoperimetric bound only up to a suboptimal multiplicative constant (by a factor $1/\ln 2$).
-Samorodnitsky [30] overcame this mismatch by proving a discrete functional inequality tailored to subsets: for functions supported on $A\subseteq\{0,1\}^{n}$ it yields the sharp Harper’s constant and characterizes equality via codimension-$k$ subcubes.
-Although Samorodnitsky’s inequality is formulated on the hypercube, recovering Harper’s optimal constant forces the right–hand side to involve the conditional average over the support set $A$.
-This quantity depends on the global geometry/size of $A$ and is not additive over coordinates; it cannot be written as an average of one-coordinate functionals.
-Consequently, the standard tensorization machinery for product spaces, which relies on coordinatewise factorization, cannot “lock onto” the correct constant.
-In this sense, even though the ambient space is product, conditioning on an arbitrary $A$ unfortunately destroys tensor structure at the level of the inequality, revealing difficulties that mirror those in genuinely non-product settings.
-In the $p$-biased setting, while a sharp edge-isoperimetric inequality for increasing sets is known (see, e.g., [15, 25]), the natural $p$-biased log-Sobolev inequalities (e.g., [7, 9]) again fail to deliver the optimal constant for indicators except in limiting regimes (such as $p\to 0$), leaving a quantitative and conceptual gap.
-These considerations highlight an urgent need to establish a Samorodnitsky-type functional inequality on the $p$-biased cube.
-A second illustration comes from mixing of censored random walks, where product structure is explicitly broken.
+In the classical uniform product setting on the hypercube $\{0,1\}^{n}$, these inequalities are well understood: sharp constants are known (e.g., the exact log-Sobolev constant [BT2006logsobolev]) and elegant proofs arise via tensorization [gross1975logarithmic], semigroup methods [Ivanisvili2024KKL, ivanisvili2020rademacher], or discrete Fourier analysis [Eldan2025Isoperimetric].
+In this paper, we depart from the product setting–by restricting to structured subsets or by working under biased measures, and aim at deriving sharp functional inequalities in these spaces and applying them to characterize the convergence behaviors of relevant random walks.
+Beyond the product setting, the study of censored random walks on increasing subsets of the hypercube is motivated by foundational challenges in Markov Chain Monte Carlo (MCMC) for constrained sampling [Jer1998, guruswami2016rapidly, MT2006, LP2017].
+A canonical example is sampling uniformly from configurations of a statistical mechanics model (e.g., critical percolation [DM2014]) that satisfy a non-local property such as a left-to-right crossing.
+While the ambient space is a product measure and trivial to sample from, the constraint set forms a highly correlated, nonproduct subset.
+Rejection sampling becomes statistically inefficient for rare events, prompting the use of local chains that remain within the constraint set.
+The natural censored dynamics re-randomizes one coordinate at a time and accepts the move only if the property is preserved.
+This process evolves on an increasing subset, and monotonicity ensures irreducibility.
+The central algorithmic question is to quantify the mixing time and the related survival/exit characteristics of this censored walk, since these quantities determine the efficiency and robustness of constrained sampling in combinatorics, statistical physics, and property testing.
+1.1 Poincaré inequality and random walks on increasing sets
+An illustration of non-product structure is motivated from censored random walks, where product structure is explicitly broken.
 If we censor the simple random walk on $\{0,1\}^{n}$ to a subset $A$ (only moves that stay in $A$ are accepted), the chain now lives on a non-product space with geometry dictated by the boundary of $A$.
-In this setting, standard product-based tools (tensorization, Fourier decompositions) no longer apply cleanly, and the behavior can vary widely: the walk may mix well, but it can also mix very poorly even when $A$ is large and connected (see [17, Examples 1–2]).
-Motivated by the question of Ding and Mossel [10, Question 1.1]—whether $t_{\mathrm{mix}}(\varepsilon)\leq O_{c,\varepsilon}(n\log n)$ holds for increasing sets of constant density—recent progress of Fei and Ferreira Pinto Jr [17] proves an $O(n^{2})$ bound by establishing a Poincaré inequality on the restricted (non-product) space.
-Their argument relies on a rather complex directed heat-flow method.
-This naturally raises a question: can simpler methods–such as an induction argument on the dimension–still be effective in such non-product spaces?
-Pursuing this direction can clarify which obstacles are intrinsic to the loss of product structure and offer a cleaner pathway toward functional inequalities in genuinely non-product settings.
-A particularly powerful paradigm that has recently emerged for proving such inequalities is the induction by restrictions combined with two-point inequalities, an approach widely adopted within the combinatorics and discrete analysis community, as exemplified by [1, 2, 3, 8, 12, 15, 16, 25, 30, 33].
-Our contributions.
-In this paper we establish two new discrete functional inequalities along this framework: one concerns a sharp $p$-biased edge-isoperimetric inequality for all real-valued increasing functions, and the other is a Poincaré inequality on increasing subsets of the hypercube.
-Both are proved by induction on the dimension combined with a two-point inequality, illustrating the versatility of this method across departures from the product setting.
-  • (1)
-    $p$-biased edge-isoperimetric inequality.
-In the uniform cube, Harper’s edge-isoperimetric theorem identifies subcubes as the unique minimizers of edge boundary, and Samorodnitsky [30] established a functional extension with sharp constants.
-For the $p$-biased hypercube endowed with product measure $\mu_{p}$, however, only weaker forms were known, typically losing sharp constants.
-We establish a new $p$-biased edge-isoperimetric inequality for real-valued increasing functions via induction on dimension and a two-point verification.
-Our result recovers the sharp $p$-biased isoperimetric inequality for increasing sets, and admits a probabilistic interpretation: increasing subcubes maximize the mean first exit time of the $p$-biased random walk among increasing sets of fixed measure.
-  • (2)
-    Poincaré inequality on increasing sets.
-Ding and Mossel [10] introduced the censored random walk on an increasing subset $A\subseteq\{0,1\}^{n}$ and asked whether it mixes in $O(n\log n)$ steps under constant density.
-They proved an $O(n^{3})$ upper bound.
-Fei and Ferreira Pinto Jr [17] established a Poincaré inequality on increasing sets that improves the spectral gap and implies an $O(n^{2})$ mixing-time upper bound.
-We give a simpler inductive proof of this Poincaré inequality, again leading to the $O(n^{2})$ mixing bound for the censored random walk.
-1.1 Functional version of the biased edge-isoperimetric inequality
-In this subsection we work on the $n$-dimensional hypercube.
-The state space is the graph $Q^{n}=(\{0,1\}^{n},E)$ with $2^{n}$ vertices indexed by binary strings, where two vertices are adjacent (written $x\sim y$) iff they differ in exactly one coordinate.
-Let $g:\{0,1\}^{n}\to\mathbb{R}$, we define the Dirichlet form of $g$ as
-$\mathcal{E}(g,g)=\underset{x\sim\mu}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}(g(x)-g(y))^{2}\right],$
-where the expectation is taken with respect to the uniform probability measure $\mu:=\mu_{1/2}^{\otimes n}$ on the hypercube.
-Furthermore, we denote by $x\sim A$ that $x$ is drawn uniformly at random from the subset $A\subseteq\{0,1\}^{n}$.
-In 2017, Samorodnitsky [30] established the following discrete functional inequality on the hypercube.
-Theorem 1.1 (Samorodnitsky [30]).
-Let $A$ be a subset of $\{0,1\}^{n}$ and let $g:\{0,1\}^{n}\to\mathbb{R}$ supported on $A$. Then
-$\mathcal{E}(g,g)\geq\frac{|A|}{2^{n-1}}\left(\underset{x\sim A}{\mathbb{E}}[|g(x)|]\right)^{2}\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right),$ (1)
-equality holds if and only if $g$ is an indicator function of a co-dimension $k$ subcube for any $0\leq k\leq n$.
-A subcube of co-dimension $k$ is a set of the form
-$\{x\in\{0,1\}^{n}:x_{i_{1}}=a_{1},x_{i_{2}}=a_{2},\dots,x_{i_{k}}=a_{k}\},$
-where $i_{1}<\cdots<i_{k}$ are coordinate indices, and $a_{1},\dots,a_{k}$ are fixed elements of $\{0,1\}$.
-The coordinates indexed by $i_{1},\dots,i_{k}$ are called the fixed coordinates; the other coordinate indices are called the active coordinates.
-Remark.
-  • (1)
-    When $g$ is an indicator function of $A$, the Samorodnitsky’s inequality (1) implies the classic edge-isoperimetric inequality on $Q^{n}$: for every $A\subset\{0,1\}^{n}$,
-    $|\nabla A|\geq|A|\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right),$ (2)
-    where the edge boundary $\nabla(A)$ is the set of edges between $A$ and its complement $A^{c}=V(Q^{n})\setminus A$.
-Equality in (2) holds if and only if $A$ is a subcube of co-dimension $k$, for any $0\leq k\leq n$.
-It is easy to see that $\mathcal{E}(\mathbbm{1}_{A},\mathbbm{1}_{A})=\frac{2|\nabla A|}{2^{n}}$.
-The fully edge isoperimetric problem for $Q^{n}$ was solved by Harper [21], Lindsey [28], Bernstein [5], and Hart [22], which states that among all subsets of $\{0,1\}^{n}$ of size $1\leq k\leq 2^{n}$, the first $k$ elements of the lexicographic ordering on $\mathcal{P}([n])$ (defined by $x<y\Leftrightarrow\max(x\Delta y)\in y$) have the smallest possible edge-boundary.
-  • (2)
-    A primary motivation for Samorodnitsky’s functional inequality is to obtain a general discrete functional inequality for real-valued functions that recovers the sharp/correct constant in (2).
-In contrast, the classical logarithmic Sobolev inequality on the hypercube only recovers (2) up to a multiplicative factor $\frac{1}{\ln 2}$:
-    $\mathcal{E}(g,g)\geq 2\cdot{\rm Ent}(g^{2}):=2\mathbb{E}[g^{2}\ln(g^{2})]-2\mathbb{E}[g^{2}]\ln(\mathbb{E}[g^{2}]).$ (3)
-  • (3)
-    To see the connection between Samorodnitsky’s inequality (1) and the log-Sobolev inequality (3), we rewrite it as
-    $\mathcal{E}(g^{\prime},g^{\prime})\geq 2\cdot\frac{|A|}{2^{n}}\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right)$
-    for nonnegative $g$, where $g^{\prime}:=g/\mathbb{E}_{A}[g]$, and $\mathbb{E}_{A}[g]$ is the conditional expectation on $A$.
-We generalize Samorodnitsky’s inequality to the $p$-biased setting so that the resulting functional inequality recovers the sharp constant in the $p$-biased edge-isoperimetric inequality.
-Let $f\in L^{2}(\{0,1\}^{n},\mu_{p}^{\otimes n})$, where $\mu_{p}$ is the Bernoulli measure on $\{0,1\}$ with $\mu_{p}(1)=p$ and $\mu_{p}(0)=1-p$.
-For $x\in\{0,1\}^{n}$, write $|x|=|\{i\in[n]:x_{i}=1\}|$, so that
-$\mu_{p}(x):=\mu_{p}^{\otimes n}(x)=p^{|x|}(1-p)^{n-|x|},\quad\mathbb{E}_{p}[f]:=\underset{x\sim\mu_{p}}{\mathbb{E}}[f(x)]=\sum\limits_{x\in\{0,1\}^{n}}f(x)\mu_{p}(x).$
-Many applications (e.g., to the study of percolation [4], threshold phenomena in random graphs [26, 31], and hardness of approximation [11]) rely upon the use of the biased measure on the discrete cube.
-The definition of Dirichlet quadratic form of $g$ with respect to the biased measure is
-$\mathcal{E}_{p}(g,g):=\underset{x\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}(g(x)-g(y))^{2}\right].$
-We consider increasing functions defined on the discrete cube $\{0,1\}^{n}$ endowed with the $p$-biased measure.
-Definition 1.2.
-A function $g:(\{0,1\}^{n},\mu_{p})\to\mathbb{R}$ is increasing if for all $x,y\in\{0,1\}^{n}$,
-$x_{i}\leq y_{i},\ \forall i\in[n]\Rightarrow g(x)\leq g(y).$
-A subset $A\subset\{0,1\}^{n}$ is called increasing if its characteristic function is increasing.
-The following ‘biased’ generalization of (2) for increasing subsets is considered folklore (see [15, 25]).
-Theorem 1.3.
-Let $A$ be an increasing subset of $\{0,1\}^{n}$ and let $0<p<1$. Then
-$p\cdot\mathcal{E}_{p}(\mathbbm{1}_{A},\mathbbm{1}_{A})\geq\mu_{p}(A)\log_{p}\mu_{p}(A),$ (4)
-equality holds in (4) if and only if $A$ is an increasing subcube.
-Inspired by (1), it is natural to seek a functional inequality on the $p$-biased cube for general real-valued (increasing) functions that, in the special case of indicator functions, reduces to the $p$-biased edge-isoperimetric inequality with the sharp constant.
-We establishes such an inequality, which recovers (4) by substituting $g=\mathbbm{1}_{A}$.
-Theorem 1.4.
-Let $A$ be a subset of $\{0,1\}^{n}$ and let $g$ be a real-valued increasing function on $\{0,1\}^{n}$ supported on $A$. Then for any $0<p<1$,
-$p\cdot\mathcal{E}_{p}(g,g)\geq\frac{\mathbb{E}_{p}[|g|]^{2}}{\mu_{p}(A)}\log_{p}\mu_{p}(A),$ (5)
-equality holds if and only if $g$ is an indicator function of an increasing subcube of co-dimension $k$ for any $0\leq k\leq n$.
-Remark.
-Note that the logarithmic Sobolev inequality (3) can be generalized to $p$-biased measure as follows, see [7, 9]:
-$\mathcal{E}_{p}(f,f)\geq\frac{1}{p(1-p)}\cdot\frac{1-2p}{\ln(1-p)-\ln(p)}\cdot{\rm Ent}_{p}(f^{2}).$ (6)
-When $f:=\mathbbm{1}_{A}$, we have the following
-$p\cdot\mathcal{E}_{p}(f,f)\geq\frac{\ln(1/p)}{1-p}\cdot\frac{1-2p}{\ln(1-p)-\ln(p)}\cdot\mu_{p}(f)\log_{p}\mu_{p}(f),$ (7)
-and hence for $p\to 0$,
-$(1+o_{p}(1))p\cdot\mathcal{E}_{p}(f,f)\geq\mu_{p}(f)\log_{p}\mu_{p}(f),$
-which again can not cover the original $p$-biased version of edge-isoperimetric inequality (4).
-We view the $p$-biased hypercube as a weighted undirected graph $Q^{n}_{p}=(\{0,1\}^{n},E,w_{p})$, where for every vertex $x\in\{0,1\}^{n}$ and coordinate index $i\in[n]$, the edge weight is defined by
-$\begin{split}w_{p}(\{x,x\oplus e_{i}\})&=p^{|x|}(1-p)^{n-|x|}+p^{|x\oplus e_{i}|}(1-p)^{n-|x\oplus e_{i}|}\\ &=p^{\sum_{j\in[n]:j\neq i}x_{j}}(1-p)^{n-\sum_{j\in[n]:j\neq i}x_{j}}\left(p^{x_{i}}(1-p)^{-x_{i}}+p^{x_{i}\oplus e_{i}}(1-p)^{-x_{i}\oplus e_{i}}\right)\\ &=p^{\sum_{j\in[n]:j\neq i}x_{j}}(1-p)^{n-\sum_{j\in[n]:j\neq i}x_{j}-1}.\end{split}$
-Thus each edge is weighted proportionally to the combined $p$-biased mass of its two endpoints.
-A random walk on a graph is a process that begins at some vertex, and at each time step moves to another vertex.
-When the graph is unweighted, the vertex the (simple) random walk moves to is chosen uniformly at random among the neighbors of the present vertex.
-When the graph is weighted, it moves to a neighbor with probability proportional to the weight of the corresponding edge.
-In [30], Samorodnitsky observed that the functional inequality (1) is equivalent to a probabilistic statement about random walks on the hypercube: subcubes maximize mean first exit time among all subsets of the same cardinality.
-Using the same argument, our new $p$-biased edge-isoperimetric inequality (5) admits an analogous probabilistic interpretation in terms of a simply random walk on the weighted hypercube $Q^{n}_{p}$.
-Let $A\subseteq\{0,1\}^{n}$ and define the following process.
-Choose a uniformly random point $x\in A$ and then moving to one of its $n$ neighbors with probability ${\rm P}_{p}(x,x\oplus e_{i})=\omega_{p}(\{x,x\oplus e_{i}\})/d_{p}(x)$, where $d_{p}(x)=\sum_{i=1}^{n}\omega_{p}(\{x,x\oplus e_{i}\})$ is the weighted degree of vertex $x$.
-Let $Y$ be the time that it takes the walk to exit $A$ for the first time.
-We refer to $\mathbb{E}_{p}[Y]$ as the mean first exit time of $A$ under the $p$-biased random walk.
-Corollary 1.5.
-Increasing subcubes maximize mean first exit time among all increasing subsets of the weighted hypercube $Q_{p}^{n}$ of the same cardinality.
-More precisely, for any increasing subset $A\subset\{0,1\}^{n}$,
-$\mathbb{E}_{p}[Y]\leq\frac{n}{\log_{p}(\mu_{p}(A))},$ (8)
-equality holds if and only if $g$ is indicator functions of increasing subcubes.
-1.2 Poincaré inequality on increasing sets
-Discrete-time random walks on finite graphs $G=(V,E)$ model stochastic processes where a walker iteratively transitions between adjacent vertices.
+Currently, there is a lack of good estimation for the mixing time of such random walks.
+Let us introduce this problem in details.
+We work on the $n$-dimensional hypercube.
+The Hamming graph is the graph $Q^{n}=(\{0,1\}^{n},E)$ with $2^{n}$ vertices indexed by binary strings, where two vertices are adjacent (written $x\sim y$) iff they differ in exactly one coordinate.
+Discrete-time random walks on the Hamming graph (or more generally, on finite graphs $G=(V,E)$) model stochastic processes where a walker iteratively transitions between adjacent vertices.
 Formally, at each timestep $t$, the walker moves from current vertex $u$ to neighbor $v$ with probability $P(u,v)=\frac{1}{{\deg}(u)}$, defining a Markov chain with state space $V$ and transition matrix $P$.
@@
 where $\|\cdot\|_{{\rm TV}}$ denotes the total variation, and $\varepsilon\in(0,1)$ governs approximation accuracy.
-Ding and Mossel [10] introduced the study of the random walk on $\{0,1\}^{n}$ censored to $A$.
+Ding and Mossel [DM2014] introduced the study of the random walk on $\{0,1\}^{n}$ censored to $A$.
 This walk, when at state $x\in A$, attempts to flip a uniformly random coordinate; it moves to the neighbor if it resides in $A$, otherwise it stays at $x$.
-Note that if we do not impose any restrictions on $A$, the censored random walk may mix well, and also may mix extremely poorly even when $A$ is large and connected, see [17, Example 1,2].
-A significant open question ([10, Question 1.1]) asked if the optimal mixing time belongs to $O_{c}(n\log n)$ holds for increasing subsets $A$ with constant density under the uniform measure.
-Question 1.6 ([10, Question 1.1]).
+Note that if we do not impose any restrictions on $A$, the censored random walk may mix well, and also may mix extremely poorly even when $A$ is large and connected, see [fei2025spectral, Example 1,2].
+A significant open question ([DM2014, Question 1.1]) asked if the optimal mixing time belongs to $O_{c}(n\log n)$ holds for increasing subsets $A$ with constant density under the uniform measure.
+Question 1 ([DM2014, Question 1.1]).
 Suppose that there exists a constant $c>0$ such that a increasing subset $A\subseteq\{0,1\}^{n}$ has measure $\mu(A)\geq c$.
@@
 When the density $\mu(A)$ is constant, Ding and Mossel established a mixing time bound of $O(n^{3})$.
-Very recently, Fei and Ferreira Pinto Jr [17] show an improved mixing time bound of $O(n^{2})$ for constant-density increasing sets via establishing the following Poincaré inequality on the non-product space.
-Theorem 1.7 ([17, Theorem 1.6]).
+Very recently, Fei and Ferreira Pinto Jr [fei2025spectral] make a breakthrough on this topic by showing an improved mixing time bound of $O(n^{2})$ for constant-density increasing sets.
+Their result is obtained via establishing the following Poincaré inequality on the non-product space.
+Theorem 1.1 ([fei2025spectral, Theorem 1.6]).
 Let $A\subseteq\{0,1\}^{n}$ be a non-empty increasing set. It holds that for all $f:A\to\mathbb{R}$,
@@
 In this paper, we provide a simple inductive proof of their Poincaré inequality, though yielding a slightly worse constant (theirs being 1 versus ours at 2).
-Specifically, we prove the following theorem, a weaker version of Theorem 1.7.
+Specifically, we use the induction method to prove the following theorem, a weaker version of Theorem 1.1.
 Theorem 1.2.
@@
 ${\rm Var}_{A}[f]\leq\frac{2}{1-\sqrt{1-\mu(A)}}\cdot\mathcal{E}_{A}(f).$ (4)
-Nevertheless, we still obtain an $O(n^{2})$ mixing time bound for increasing sets $A$ of constant density via standard Markov chain theory ([27, Theorem 12.4]).
-Corollary 1.9.
+Nevertheless, we still obtain an $O(n^{2})$ mixing time bound for increasing sets $A$ of constant density via standard Markov chain theory ([LP2017, Theorem 12.4]).
+Theorem 1.3 ([fei2025spectral]).
 Let $A\subseteq\{0,1\}^{n}$ be a non-empty increasing set. Then the random walk on $\{0,1\}^{n}$ censored to $A$ has mixing time
 $t_{{\rm mix}}\leq\frac{2n}{\mu(A)}\cdot\log(4\cdot 2^{n}\mu(A)).$
-Organization.
+Fei and Ferreira Pinto Jr.
+[fei2025spectral] develop an $L^{2}$ functional–analytic framework for increasing sets $A\subseteq\{0,1\}^{n}$.
+They introduce a directed gradient that counts only order-consistent edges (down moves), which defines a directed Laplacian $L_{A}^{\to}$ and energy $\mathcal{E}_{A}^{\to}(f)$.
+They then run the monotone heat flow generated by $L_{A}^{\to}$ and prove an energy–variance dissipation inequality along the flow, yielding a directed Poincaré inequality with constant $C(\mu(A))=O(1)$ whenever $\mu(A)$ is bounded away from 0.
+A second ingredient is an approximate FKG inequality on increasing sets that supplies quantitative positive-correlation bounds.
+Together these tools give an optimal (up to constants) spectral gap for the censored walk on $A$.
+In contrast, our method is designed to be simple and easy to check.
+We establish a slightly weaker, yet asymptotically optimal, Poincaré inequality by a classical induction-on-dimension argument.
+The proof restricts functions to lower-dimensional subcubes and uses the inherent increasing structure $A_{0}\subseteq A_{1}$ to decompose the variance and the Dirichlet form fiberwise, ultimately reducing the task to an explicit five-point inequality that can be verified directly.
+The resulting constant is within a factor $2$ of the optimal value and already implies the same mixing bound $O(n^{2})$ for censored walks.
+Because every step is combinatorial and closed-form, the proof is short, self-contained, and readily adaptable.
+1.2 Biased Samorodnitsky’s inequality and exit time of random walks from increasing sets
+Another illustration of non-product structure is motivated from characterizing the mean first exit time of a random walk from an increasing set.
+Let $g:\{0,1\}^{n}\to\mathbb{R}$, we define the Dirichlet form of $g$ as
+$\mathcal{E}(g,g)=\underset{x\sim\mu}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}(g(x)-g(y))^{2}\right],$
+where the expectation is taken with respect to the uniform probability measure $\mu:=\mu_{1/2}^{\otimes n}$ on the hypercube.
+Furthermore, we denote by $x\sim A$ that $x$ is drawn uniformly at random from the subset $A\subseteq\{0,1\}^{n}$.
+In 2017, Samorodnitsky [Samorodnitsky17] established the following discrete functional inequality on the hypercube.
+Theorem 1.4 (Samorodnitsky [Samorodnitsky17]).
+Let $A$ be a subset of $\{0,1\}^{n}$ and let $g:\{0,1\}^{n}\to\mathbb{R}$ supported on $A$. Then
+$\mathcal{E}(g,g)\geq\frac{|A|}{2^{n-1}}\left(\underset{x\sim A}{\mathbb{E}}[|g(x)|]\right)^{2}\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right),$ (5)
+equality holds if and only if $g$ is an indicator function of a co-dimension $k$ subcube for any $0\leq k\leq n$.
+When $g$ is an indicator function of $A$, Samorodnitsky’s inequality Eq. 5 implies the classic edge-isoperimetric inequality on $Q^{n}$: for every $A\subset\{0,1\}^{n}$,
+$|\nabla A|\geq|A|\cdot\log_{2}\left(\frac{2^{n}}{|A|}\right),$ (6)
+where the edge boundary $\nabla(A)$ is the set of edges between $A$ and its complement $A^{c}=V(Q^{n})\setminus A$.
+Equality in Eq.
+6 holds if and only if $A$ is a subcube of co-dimension $k$, for any $0\leq k\leq n$.
+Besides the motivation from generalization of Haper’s edge-isoperimetric inequality, Samorodnitsky’s inequality admits a natural interpretation in languages of random walk on the hypercube.
+Specifically, it is equivalent to a sharp estimate of mean first exit time of a random walk from a subset with a given size [Samorodnitsky17].
+We now generalize Samorodnitsky’s inequality to the $p$-biased setting so that the resulting functional inequality recovers the sharp constant in the $p$-biased edge-isoperimetric inequality.
+Let $f\in L^{2}(\{0,1\}^{n},\mu_{p}^{\otimes n})$, where $\mu_{p}$ is the Bernoulli measure on $\{0,1\}$ with $\mu_{p}(1)=p$ and $\mu_{p}(0)=1-p$.
+For $x\in\{0,1\}^{n}$, write $|x|=|\{i\in[n]:x_{i}=1\}|$, so that
+$\mu_{p}(x):=\mu_{p}^{\otimes n}(x)=p^{|x|}(1-p)^{n-|x|},\quad\mathbb{E}_{p}[f]:=\underset{x\sim\mu_{p}}{\mathbb{E}}[f(x)]=\sum\limits_{x\in\{0,1\}^{n}}f(x)\mu_{p}(x).$
+Many applications (e.g., to the study of percolation [BKS1999], threshold phenomena in random graphs [KLMM2024globalhyper2024, Talagrand1994randomgraph], and hardness of approximation [DS2005Annal]) rely upon the use of the biased measure on the discrete cube.
+The definition of Dirichlet quadratic form of $g$ with respect to the biased measure is
+$\mathcal{E}_{p}(g,g):=\underset{x\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}(g(x)-g(y))^{2}\right].$
+We consider increasing functions defined on the discrete cube $\{0,1\}^{n}$ endowed with the $p$-biased measure.
+Definition 1.
+A function $g:(\{0,1\}^{n},\mu_{p})\to\mathbb{R}$ is increasing if for all $x,y\in\{0,1\}^{n}$,
+$x_{i}\leq y_{i},\ \forall i\in[n]\Rightarrow g(x)\leq g(y).$
+A subset $A\subset\{0,1\}^{n}$ is called increasing if its characteristic function is increasing.
+The following ‘biased’ generalization of Harper’s edge-isoperimetric inequality for increasing subsets is considered folklore (see [EKL2019biased, Kahnkalai07]).
+Theorem 1.5.
+Let $A$ be an increasing subset of $\{0,1\}^{n}$ and let $0<p<1$. Then
+$p\cdot\mathcal{E}_{p}(\mathbbm{1}_{A},\mathbbm{1}_{A})\geq\mu_{p}(A)\log_{p}\mu_{p}(A),$ (7)
+equality holds in Eq. 7 if and only if $A$ is an increasing subcube.
+Inspired by Eq.
+5, it is natural to seek a functional inequality on the $p$-biased cube for general real-valued (increasing) functions that, in the special case of indicator functions, reduces to the $p$-biased edge-isoperimetric inequality with the sharp constant.
+We establishes such an inequality, which recovers Eq.
+7 by substituting $g=\mathbbm{1}_{A}$.
+Theorem 1.6.
+Let $A$ be a subset of $\{0,1\}^{n}$ and let $g$ be a real-valued increasing function on $\{0,1\}^{n}$ with support $A$. Then for any $0<p<1$,
+$p\cdot\mathcal{E}_{p}(g,g)\geq\frac{\mathbb{E}_{p}[|g|]^{2}}{\mu_{p}(A)}\log_{p}\mu_{p}(A),$ (8)
+equality holds if and only if $g$ is an indicator function of an increasing subcube of co-dimension $k$ for any $0\leq k\leq n$.
+We view the $p$-biased hypercube as a weighted undirected graph $Q^{n}_{p}$, where for every vertex $x\in\{0,1\}^{n}$ and coordinate index $i\in[n]$, the edge weight is defined by
+$\begin{split}w_{p}(\{x,x\oplus e_{i}\})&=\mu_{p}^{\otimes n}(x)+\mu_{p}^{\otimes n}(x\oplus e_{i})=\mu_{p}^{\otimes(n-1)}(x_{\backslash i}),\end{split}$
+where $x_{\backslash i}=(x_{1},...,x_{i-1},x_{i+1},...,x_{n})$. Thus each edge is weighted proportionally to the combined $p$-biased mass of its two endpoints.
+A random walk on a graph is a process that begins at some vertex, and at each time step moves to another vertex.
+When the graph is unweighted (i.e., $p=1/2$), the vertex the (simple) random walk moves to is chosen uniformly at random among the neighbors of the present vertex.
+When the graph is weighted, it moves to a neighbor with probability proportional to the weight of the corresponding edge.
+In [Samorodnitsky17], Samorodnitsky observed that the functional inequality Eq.
+5 is equivalent to a probabilistic statement about random walks on the hypercube: subcubes maximize mean first exit time among all subsets of the same cardinality.
+Using the same argument, our new $p$-biased edge-isoperimetric inequality Eq.
+8 admits an analogous probabilistic interpretation in terms of a simply random walk on the weighted hypercube $Q^{n}_{p}$.
+Let $A\subseteq\{0,1\}^{n}$ and define the following random walk $X_{t}$.
+The random walk starts at a random point $X_{0}=x\in A$ following from the $p$-biased conditional law, i.e., $\nu(x):=\frac{\mu_{p}(x)}{\mu_{p}(A)},\forall x\in A$, and at each step, moves to one of its $n$ neighbors with probability ${\rm P}_{p}(x,x\oplus e_{i})=\omega_{p}(\{x,x\oplus e_{i}\})/d_{p}(x)$, where $d_{p}(x)=\sum_{i=1}^{n}\omega_{p}(\{x,x\oplus e_{i}\})$ is the weighted degree of vertex $x$.
+Let $Y$ be the (random) time that it takes the walk to exit $A$ for the first time for the $p$-biased walk.
+We refer to $T(A,p)=\mathbb{E}[Y]$ as the mean first exit time of $A$ under the $p$-biased walk.
+We further introduce the weighted mean first exit time as
+$\tilde{T}(A,p):=\mathbb{E}\left[\sum\limits_{t=0}^{Y-1}\phi(X_{t})\right],{\rm where}\ \phi(x):=\frac{\mu_{p}(x)}{d_{p}(x)}.$
+Theorem 1.7.
+Let $A\subset\{0,1\}^{n}$ be increasing and $0<p<1$. Then
+  • (1)
+    $\tilde{T}(A,p)\leq\frac{p}{\log_{p}\mu_{p}(A)}$ with equality if and only if $A$ is an increasing subcube.
+  • (2)
+    $T(A,p)\leq\frac{p}{\min\{p,1-p\}}\cdot\frac{n}{\log_{p}(\mu_{p}(A))}$ with equality if and only if $p=\tfrac{1}{2}$ and $A$ is an increasing subcube.
+Example 1 (Mean first exit time for the $p$-biased walk on increasing subcubes).
+Fix $0<p<1$ and $n\in\mathbb{N}$. Let $C:=\{x\in\{0,1\}^{n}:x_{1}=\cdots=x_{k}=1\}$ be the increasing subcube of codimension $k$, i.e., $k=\log_{p}\mu_{p}(A)$. Then
+$T(C,p)=1+2p\left(\frac{n}{k}-1\right).$
+In particular, for $p=\tfrac{1}{2}$ this reduces to $T(C,p)=n/k$.
+1.3 Organization
 This paper is organized as follows.
-Section 2 is devoted to our first main result: a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions.
+Section 2 presents our first main result: an inductive proof of a Poincaré inequality on increasing subsets of the cube, which yields an $O(n^{2})$ upper bound on the mixing time of the censored random walk.
+The arguments follow the induction + five-point paradigm, though the analytic verification is much simple.
+Section 3 turns to our second main result: a sharp $p$-biased edge-isoperimetric inequality for real-valued increasing functions.
 We develop the induction-on-dimension framework and verify the required two-point inequality in this setting.
-Section 3 turns to our second main result: an inductive proof of a Poincaré inequality on increasing subsets of the cube, which yields an $O(n^{2})$ upper bound on the mixing time of the censored random walk.
-The arguments once again follow the induction + five-point paradigm, though the analytic verification is much simple than the biased case.
-Finally, Section 4 presents concluding remarks and open problems.
-We highlight two directions for future research: (i) developing analytic proofs of weaker forms of Harper’s edge-isoperimetric inequality, and (ii) establishing suitable (possibly modified) log-Sobolev inequalities on increasing sets, which are closely related to the Ding–Mossel conjecture.
+Section 4 explains how inequality Eq.
+8 in Theorem 1.6 leads to bounds on the mean survival (first-exit) time of random walks from increasing sets (Theorem 1.7).
+Section 5 verifies the mean survival time of random walks from increasing subcubes (1).
+Finally, Section 6 highlight a direction for future work: establishing suitable (possibly modified) log-Sobolev inequalities on increasing sets, which are closely related to the Ding–Mossel conjecture.
+2 Proof of Theorem 1.2
+2.1 Preliminaries
+In this subsection, we outline some notation and terminology, and present some simple lemmas for inductive proof of Theorem 1.2.
+In the language of Boolean functions, there are two different ways to prove inequalities about $f:\{0,1\}^{n}\to\mathbb{R}$ by induction on $n$.
+One method, which might be called “induction by derivatives”, decomposes a function $f$ as follows: $f(x)=(-1)^{x_{i}}g_{1}(y)+g_{0}(y)$ for some functions $g_{0},g_{1}$ on the (n1)(n-1)-dimensional cube, where $y=(x_{1},\dots,x_{n-1})$.
+Such a decomposition is motivated from Fourier expansion.
+This induction method underlies classic proofs of Bonami’s lemma and the hypercontractive inequality on the hypercube [Ryanbook2014].
+The other method, which we term “induction by restrictions”, proceeds via the restricted functions $g_{a}$ obtained by fixing the last coordinate to $a\in\{0,1\}$.
+One then expresses Dirichlet forms (or total influence) and variances recursively in terms of $g_{0}$ and $g_{1}$, applies the inductive hypothesis in dimension n1n-1, and controls the remaining cross-terms by verifying an explicit two-point (and occasionally four/five-point) inequality.
+We prove both results within the standard induction-by-restrictions, widely used in combinatorics and discrete analysis (see, e.g., [alpay2025lowerboundsdyadicsquare, beltran2023sharpisoperimetricinequalitieshypercube, BIMP2025young, CTKS2025inequalities, durcik2024sharpisoperimetricinequalitieshamming, EKL2019biased, FS2007CPCKKL, Kahnkalai07, Samorodnitsky17, yu2025averagedistancelevel1fourier]).
+Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V^{n}$.
+Let $V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$ and $A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}$.
+Let $a_{0}:=\mu^{n-1}(A_{0}),a_{1}:=\mu^{n-1}(A_{1})$, and note that
+$\begin{split}a:=\mu^{n}(A)=\frac{1}{2}\mu^{n-1}(A_{1})+\frac{1}{2}\mu^{n-1}(A_{0})=\frac{a_{0}+a_{1}}{2}.\end{split}$ (9)
+Define $f_{0}:A_{0}\to\mathbb{R},\ f_{1}:A_{1}\to\mathbb{R}$ by
+$f_{0}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},0)\quad\text{and}\quad f_{1}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},1).$
+By expanding the Dirichlet form with respect to the conditional expectation on the $n$th variable, we can readily express the Dirichlet form of the original function in terms of those of its restricted versions.
+Lemma 2.1.
+For any $f:A\to\mathbb{R}$, we have
+$\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$ (10)
+Proof.
+We use $x^{\prime}$ to denote the first (n1)(n-1)-coordinates of $x$. First, separate the contribution of the last coordinate from that of the first n1n-1 coordinates; this gives
+$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\sum\limits_{i=1}^{n-1}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right]\\ &+\frac{1}{4}\cdot\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus n}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1-x_{n})\in A\}}\right]\end{split}$
+Next, write $x=(x^{\prime},x_{n})$ with $x^{\prime}\in\{0,1\}^{n-1}$ and condition on $x_{n}\in\{0,1\}$. Then
+$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},1)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1)\in A\}}\right]\end{split}$
+Now split according to whether the flipped point stays in the same fiber $A_{b}$ for $b\in\{0,1\}$ or moves across fibers; this becomes
+$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{1}\}}\right]\\ &+\frac{a_{0}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{0}\}}\right]\\ &+\frac{|A_{0}|}{4|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right]+\frac{|A_{0}|}{4|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right],\end{split}$
+where in the second equality, we observe that
+$\begin{split}\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}&\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\right]=\frac{1}{|A|}\sum\limits_{x^{\prime}\in A_{1}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\\ &=\frac{1}{|A|}\sum\limits_{x^{\prime}\in A_{0}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}=\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right].\end{split}$
+Similarly, by conditioning on the $n$th coordinate, the variance of the original function can be expressed in terms of the variances of its restricted functions.
+Lemma 2.2.
+For any $f:A\to\mathbb{R}$, we have
+$\begin{split}{\rm Var}_{A}[f]=&\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]\\ &+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.\end{split}$ (11)
+Proof.
+Note that
+$\begin{split}{\rm Var}_{A}[f]=\underset{x\sim A}{\mathbb{E}}[|f-\underset{x\sim A}{\mathbb{E}}f|^{2}]=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}].\end{split}$
+Let $\mu=\underset{x\sim A}{\mathbb{E}}[f(x)],\mu_{1}=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]$ and $\mu_{0}=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]$. Then
+$\mu=\frac{1}{|A|}\sum_{x\in A}f(x)=\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]+\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]=\frac{a_{1}}{2a}\mu_{1}+\frac{a_{0}}{2a}\mu_{0}.$ (12)
+Thus by the formula $\mathbb{E}[(f-b)^{2}]={\rm Var}[f]+(\mathbb{E}[f]-b)^{2}$, we have that
+$\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]={\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2},$
+and
+$\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]={\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}.$
+Thus we have
+$\begin{split}{\rm Var}_{A}[f]&=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{1}a_{0}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(\mu_{1}-\mu_{0})^{2}\end{split}$
+2.2 Induction part
+Our goal is to prove that for every function $f:A\to\mathbb{R}$,
+$\mathcal{E}_{A}(f)\geq c\cdot(1-\sqrt{1-a})\cdot{\rm Var}_{A}[f],$ (13)
+where $c>0$ is a universal constant (we may take $c=0.5$).
+It is easy to verify that the theorem holds for $n=1$.
+Let $n\geq 2$, and suppose the statement of the theorem holds when $n$ is replaced by n1n-1.
+Applying Lemma 2.1 to split the Dirichlet form of the original function into two restricted functions:
+$\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$
+By inductive assumption, we have
+$\begin{split}&\mathcal{E}_{A_{1}}(f_{1})\geq c\cdot(1-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}],\\ &\mathcal{E}_{A_{0}}(f_{0})\geq c\cdot(1-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}].\end{split}$ (14)
+Thus our goal is to show that
+$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\geq\frac{a_{1}}{2a}(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]\\ &+\frac{a_{0}}{4a}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}(c-c\sqrt{1-a}){\rm Var}_{A}[f].\end{split}$ (15)
+We also use Lemma 2.2 to split the variance of the original function to two restricted functions:
+${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$
+Thus it suffices to show that
+$\begin{split}&\frac{a_{1}(c-c\sqrt{1-a_{1}})}{2a}{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}(c-c\sqrt{1-a_{0}})}{2a}{\rm Var}_{A_{0}}[f_{0}]\\ &\quad\ +\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}\frac{a_{1}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{0}}[f_{0}]\\ &\quad\ +\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}(c-c\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (16)
+It is equivalent to show that
+$\begin{split}&ca_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+ca_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]\\ &+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\underset{?}{\geq}c\frac{a_{1}a_{0}}{a_{0}+a_{1}}(1-\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (17)
+Lemma 2.3.
+We may assume both $f_{0},f_{1}$ to be constant on $A_{0}$ and on $A_{1}\setminus A_{0}$.
+Proof.
+Let $\mu_{1}:=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]$, $\alpha:=\mathbb{E}_{x^{\prime}\sim A_{0}}[f_{1}]$, $\beta:=\mathbb{E}_{x^{\prime}\sim A_{1}\setminus A_{0}}[f_{1}]$ and $\gamma:=\mathbb{E}_{x^{\prime}\sim A_{0}}[f_{0}]$.
+Note that
+$\begin{split}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}f_{1}|^{2}]&=\frac{|A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}\\ &+\frac{|A_{1}\setminus A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{1}\setminus A_{0}|}\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}\\ &=\frac{a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}],\end{split}$ (18)
+and
+$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}(x^{\prime})-\mu_{0}-(f_{1}(x^{\prime})-\mu_{1})+\mu_{0}-\mu_{1}|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}-\mu_{0}|^{2}\right]+\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{1}-\mu_{1}|^{2}\right]+2\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &+(\mu_{0}-\mu_{1})^{2}+2(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (19)
+For the sake of simplicity, we denote $C_{1}=\frac{1}{2}\cdot a_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}})$ and $C_{0}=\frac{1}{2}\cdot a_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}})$. Then
+$\begin{split}&C_{1}\cdot{\rm Var}_{A_{1}}[f_{1}]+C_{0}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+C_{1}\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]\\ &+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &+\frac{a_{0}}{2}(\mu_{0}-\mu_{1})^{2}+a_{0}(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (20)
+Now we focus on the term
+$\begin{split}&\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]\\ &\quad+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &=\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)|\mu_{1}-f_{1}(x^{\prime})|^{2}+\left(C_{0}+\frac{a_{0}}{2}\right)|f_{0}(x^{\prime})-\mu_{0}|^{2}\\ &\quad+a_{0}(f_{0}(x^{\prime})-\mu_{0})(\mu_{1}-f_{1}(x^{\prime}))\\ &=\frac{1}{|A_{0}|}\sum_{x^{\prime}\in A_{0}}\vec{b}_{x^{\prime}}^{\top}G_{x^{\prime}}\vec{b}_{x^{\prime}},\end{split}$ (21)
+where
+$\vec{b}_{x^{\prime}}=\begin{bmatrix}f_{0}(x^{\prime})-\mu_{0}\\ \mu_{1}-f_{1}(x^{\prime})\end{bmatrix},\\ G_{x^{\prime}}=\begin{bmatrix}a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}&\frac{a_{0}}{2}\\ \frac{a_{0}}{2}&a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}\end{bmatrix}$.\\
+We claim that the matrix $G_{x^{\prime}}$ is positive semidefinite.
+Note that $G_{x^{\prime}}$ is a symmetric matrix, then we shall use the basic fact that a symmetric matrix $G=\begin{bmatrix}b&c\\ c&d\end{bmatrix}\in\mathbb{R}^{2\times 2}$ is positive semidefinite if and only if $b\geq 0,d\geq 0$ and $bd-c^{2}\geq 0$.
+It is easy to see that $a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}\geq 0$ and $a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}\geq 0$ as $a_{0}\leq a\leq a_{1}$. Thus it suffices to verify
+$\frac{a_{0}^{2}}{4}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-\frac{a_{0}^{2}}{4}\geq 0.$ (22)
+Let $s:=\sqrt{1-a_{0}}$ and $t:=\sqrt{1-a_{1}}$, then $0\leq t\leq s\leq 1$, $1-a=(s^{2}+t^{2})/2$ and
+$\begin{split}&(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-1\\ &=\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-s+1\right)\cdot\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-t+1\right)-1\\ &=-\frac{1}{2}\cdot(s+t-2)\cdot(\sqrt{2}\sqrt{s^{2}+t^{2}}-s-t)\geq 0,\end{split}$ (23)
+where the final one is obviously nonnegative as $s+t-2\leq 0$ and $2(s^{2}+t^{2})\geq(s+t)^{2}$.
+By the second-order condition of convexity, we know that for any $x^{\prime}\in A_{0}$, the quadratic function $G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))$ is a convex function.
+Thus by Jensen’s inequality, we have
+$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))]\\ &\geq G\left(\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})-\mu_{0}],\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[\mu_{1}-f_{1}(x^{\prime})]\right)=G(0,\mu_{1}-\alpha),\end{split}$
+and
+$\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]\geq\left(\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[f_{1}-\mu_{1}]\right)^{2}=(\beta-\mu_{1})^{2}.$
+Hence replacing $f_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of Eq. 17. ∎
+2.3 Five-point Inequality Part
+Now let $\alpha$ be the value of $f_{1}$ on $A_{0}$ and $\beta$ be the value of $f_{1}$ on $A_{1}\setminus A_{0}$. Let $\gamma$ be the value of $f_{0}$ on $A_{0}$. Then we compute
+$\begin{split}&{\rm Var}_{A_{1}}[f_{1}]=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}^{2}]-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}]^{2}\\ &=\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta^{2})\\ &-\left(\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta)\right)^{2}\\ &=\frac{1}{a_{1}}(a_{0}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot\beta^{2})-\frac{1}{a_{1}^{2}}(a_{0}\cdot\alpha+(a_{1}-a_{0})\cdot\beta)^{2}\\ &=\frac{a_{1}a_{0}-a_{0}^{2}}{a_{1}^{2}}\cdot(\alpha-\beta)^{2},\end{split}$
+and
+${\rm Var}_{A_{0}}[f_{0}]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}^{2}]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}]^{2}=\frac{a_{0}\cdot 2^{n-1}\cdot\gamma^{2}}{2^{n-1}\cdot a_{0}}-\left(\frac{a_{0}\cdot 2^{n-1}\cdot\gamma}{2^{n-1}\cdot a_{0}}\right)^{2}=0.$
+Thus to verify Eq. 17, it suffices to prove the following five-point inequality with $c=\frac{1}{2}$.
+$\begin{split}&c(\sqrt{1-a}-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})\cdot(\alpha-\beta)^{2}+\frac{a_{1}}{2}\cdot(\gamma-\alpha)^{2}\\ &\geq\frac{c}{a_{1}+a_{0}}\cdot(1-\sqrt{1-a})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}.\end{split}$ (24)
+Note that
+$\sqrt{1-a}-\sqrt{1-a_{1}}=\frac{1-a-(1-a_{1})}{\sqrt{1-a}+\sqrt{1-a_{1}}}=\frac{a_{1}-a_{0}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})},$
+and
+$1-\sqrt{1-a}=\frac{1-(1-a)}{1+\sqrt{1-a}}=\frac{a_{1}+a_{0}}{2(1+\sqrt{1-a})}.$
+Then it suffice to prove that
+$\begin{split}&\frac{c\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})}+\frac{a_{1}}{2}\cdot(\alpha-\gamma)^{2}\\ &\geq\frac{c}{2(1+\sqrt{1-a})}\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\\ &\Longleftrightarrow c\cdot(1+\sqrt{1-a})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}\\ &\quad\quad\ +a_{1}\cdot(1+\sqrt{1-a})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &\quad\quad\ -c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\geq 0.\end{split}$ (25)
+Note that $a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)=(a_{0}-a_{1})\cdot(\alpha-\beta)+a_{1}\cdot(\alpha-\gamma).$ After some simple calculations, the inequality to be verified becomes
+$\begin{split}&c(1-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}\\ &+a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\cdot(a_{0}-a_{1})\cdot(\alpha-\beta)\cdot(\alpha-\gamma)\geq 0.\end{split}$ (26)
+If $\alpha=\gamma$, the inequality is trivially satisfied.
+Thus, assume $\alpha\neq\gamma$ and set $T:=\frac{\alpha-\beta}{\alpha-\gamma}\cdot(a_{1}-a_{0})$.
+Then the above inequality can be rewritten as a quadratic in $T$:
+$\begin{split}&c\cdot(1-\sqrt{1-a_{1}})\cdot T^{2}+2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot T\\ &a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\geq 0.\end{split}$ (27)
+We may rewrite the left-hand side as a quadratic polynomial in $T$ ($AT^{2}+BT+C$), whose coefficients depend only on $a_{1},a_{0}$:
+$\begin{split}&A:=c\cdot(1-\sqrt{1-a_{1}})\geq 0\\ &B:=2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\\ &C:=a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}}).\end{split}$
+To ensure the quadratic is nonnegative for all $T$, it suffices to check that its discriminant satisfies $\Delta=B^{2}-4AC\leq 0$ with $c=0.5$, which will conclude the proof.
+Lemma 2.4.
+$B^{2}-4AC\leq 0$.
+Proof.
+Denote $x=a_{0},y=a_{1}$. Simplifying and rearranging, we need to prove that for any $0\leq x\leq y\leq 1$, we have
+$\begin{split}&y\cdot\left(\sqrt{1-\frac{x+y}{2}}+\sqrt{1-y}\right)-2(1-\sqrt{1-y})\cdot\left(1+\sqrt{1-\frac{x+y}{2}}-\frac{y}{2}\right)\leq 0\\ &\Leftrightarrow\left(\sqrt{1-\frac{x+y}{2}}+1\right)\cdot(y-2+2\sqrt{1-y})\leq 0.\end{split}$
+It suffices to show that $(y-2+2\sqrt{1-y})\leq 0$. It is trivially true as
+$\begin{split}&y-2+2\sqrt{1-y}\leq 0\\ &\Leftrightarrow 2\sqrt{1-y}\leq 2-y\\ &\Leftrightarrow y^{2}\geq 0.\end{split}$
 3 Proof of Theorem 1.6
@@
 $\left\langle f,g\right\rangle_{p}=\underset{x\sim\mu_{p}}{\mathbb{E}}[f(x)g(x)]=\int_{\{0,1\}^{n}}f(x)g(x)\mu_{p}(x)=\sum\limits_{x\in\{0,1\}^{n}}f(x)\cdot g(x)\mu_{p}(x).$
-Definition 2.1.
+Definition 2.
 For $q\geq 1$, the $q$-norm w.r.t. $p$-biased measure of $f:\{0,1\}^{n}\to\mathbb{R}$ is
 $\|f\|_{q,\mu_{p}}:=\left(\underset{x\sim\mu_{p}}{\mathbb{E}}[|f(x)|^{q}]\right)^{1/q}.$
-In the language of Boolean functions, there are two different ways to prove inequalities about $f:\{0,1\}^{n}\to\mathbb{R}$ by induction on $n$.
-One method, which might be called “induction by derivatives”, decomposes a function $f$ as follows: $f(x)=(-1)^{x_{i}}g_{1}(y)+g_{0}(y)$ for some functions $g_{0},g_{1}$ on the (n1)(n-1)-dimensional cube, where $y=(x_{1},\dots,x_{n-1})$.
-Such a decomposition is motivated from Fourier expansion.
-This induction method underlies classic proofs of Bonami’s lemma and the hypercontractive inequality on the hypercube [29].
-The other method, which we term “induction by restrictions”, proceeds via the restricted functions $g_{a}$ obtained by fixing the last coordinate to $a\in\{0,1\}$.
-One then expresses Dirichlet forms (or total influence) and variances recursively in terms of $g_{0}$ and $g_{1}$, applies the inductive hypothesis in dimension n1n-1, and controls the remaining cross-terms by verifying an explicit two-point (and occasionally four/five-point) inequality.
-This strategy—used in our proofs of Theorems 1.4 and 1.8—systematically reduces high-dimensional functional inequalities to finite-dimensional checks and often yields sharp constants.
-Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V$. Let
-$V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$
-and
-$A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}.$
-Let $\mu_{p}^{n-1}(A_{0})=a_{0},\mu_{p}^{n-1}(A_{1})=a_{1}$, and $\mu_{p}^{n}(A)=a=pa_{1}+(1-p)a_{0}$. Define $g_{0},g_{1}:(\{0,1\}^{n-1},\mu_{p}^{n-1})\to\mathbb{R}$ by
-$g_{0}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},0)\quad\text{and}\quad g_{1}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},1),$
-and the support of $g_{1}$ and $g_{0}$ are the whole set $A_{1}$ and $A_{0}$.
+Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V$.
+Let $V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$ and $A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}.$
+Let $\mu_{p}^{n-1}(A_{0})=a_{0},\mu_{p}^{n-1}(A_{1})=a_{1}$, and $\mu_{p}^{n}(A)=a=pa_{1}+(1-p)a_{0}$.
+Define $g_{0},g_{1}:(\{0,1\}^{n-1},\mu_{p}^{n-1})\to\mathbb{R}$ by $g_{0}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},0)$ and $g_{1}(x_{1},\dots,x_{n-1}):=g(x_{1},\dots,x_{n-1},1)$ and the support of $g_{1}$ and $g_{0}$ are the whole set $A_{1}$ and $A_{0}$.
 Note that if $g$ is an increasing function, then the set $A$ is also an increasing set, since it is the support of an increasing function $g$.
 This implies $A_{0}\subset A_{1}$.
-Fact 2.2.
+Fact 3.1.
 If $A$ is an increasing set, then $A_{0}\subset A_{1}$.
@@
 Note that
-$\begin{split}\mathcal{E}_{p}^{n}(g,g)&=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\underset{x_{n}\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},x_{n})}(g(x^{\prime},x_{n})-g(y))^{2}\right]\\ &=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[p\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}+(1-p)\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}\right]\\ &=p\cdot\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in V_{1}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}+\sum\limits_{y\in V_{0}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}\right]\\ &+(1-p)\cdot\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in V_{1}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}+\sum\limits_{y\in V_{0}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}\right]\\ &=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+p\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\\ &+(1-p)\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}.\end{split}$
+$\begin{split}\mathcal{E}_{p}^{n}(g,g)&=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\underset{x_{n}\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},x_{n})}(g(x^{\prime},x_{n})-g(y))^{2}\right]\\ &=p\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},1)}(g(x^{\prime},1)-g(y))^{2}\right]\\ &+(1-p)\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim(x^{\prime},0)}(g(x^{\prime},0)-g(y))^{2}\right].\end{split}$
+Then we decompose each sum over $V_{1}$ and $V_{0}$ to obtain
+$\begin{split}\mathcal{E}_{p}^{n}(g,g)&=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+p\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\\ &+(1-p)\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &=p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})+(1-p)\mathcal{E}_{p}^{n-1}(g_{0},g_{0})+\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}.\end{split}$
 We also need to use some simple properties of the function $f(t):=\frac{\log_{p}t}{t}$.
@@
 We shall make several simple assumptions:
-  • We may assume without loss of generality that $g\geq 0$, since replacing $g$ by its absolute value preserves the right-hand side of (5), while it only decreases the left-hand side:
+  • We may assume without loss of generality that $g\geq 0$, since replacing $g$ by its absolute value preserves the right-hand side of Eq. 8, while it only decreases the left-hand side:
     $\begin{split}\mathcal{E}_{p}(g,g)&=\underset{x\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}|g(x)-g(y)|^{2}\right]\\ &\geq\underset{x\sim\mu_{p}}{\mathbb{E}}\left[\sum\limits_{y\in\{0,1\}^{n}:y\sim x}(|g(x)|-|g(y)|)^{2}\right]=\mathcal{E}_{p}(|g|,|g|).\end{split}$
@@
 Then $\mathbb{E}_{p}[g]^{2}=\left(\sum_{x\in B}g(x)\mu_{p}(x)\right)^{2}$.
-Hence, replacing the support of $g$ with the whole set $A$ only increases the right-hand side of (5).
+Hence, replacing the support of $g$ with the whole set $A$ only increases the right-hand side of Eq.
+8.
 It is easy to check that the theorem holds for $n=1$.
@@
 By inductive assumption, we have
-$\begin{split}&p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})\geq\frac{\mathbb{E}_{p}[g_{1}]^{2}}{a_{1}}\log_{p}a_{1}\\ &p\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\geq\frac{\mathbb{E}_{p}[g_{0}]^{2}}{a_{0}}\log_{p}a_{0}.\end{split}$ (15)
+$p\mathcal{E}_{p}^{n-1}(g_{1},g_{1})\geq\frac{\mathbb{E}_{p}[g_{1}]^{2}}{a_{1}}\log_{p}a_{1},\quad p\mathcal{E}_{p}^{n-1}(g_{0},g_{0})\geq\frac{\mathbb{E}_{p}[g_{0}]^{2}}{a_{0}}\log_{p}a_{0}.$ (30)
 Thus our goal is to show that
@@
     where the first inequality is due to $a_{j}=\mu_{p}^{n-1}(A_{j})\leq 1$ and the second inequality is due to Jensen’s inequality.
-Hence replacing $g_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of (16).
+Hence replacing $g_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of Eq.
+31.
 Now let $\alpha$ be the value of $g_{1}$ on $A_{0}$ and $\gamma$ be the value of $g_{1}$ on $A_{1}\setminus A_{0}$. Let $\beta$ be the value of $g_{0}$ on $A_{0}$. Note that
-$\begin{split}&s_{1}:=\mathbb{E}_{p}[g_{1}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)=\alpha a_{0}+\gamma(a_{1}-a_{0}),\\ &s_{0}:=\mathbb{E}_{p}[g_{0}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)=\beta a_{0},\\ &\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &\quad\ =\sum\limits_{x^{\prime}\in A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})+\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})\\ &\quad\ =(\alpha-\beta)^{2}\cdot a_{0}+(a_{1}-a_{0})\cdot\gamma^{2}.\end{split}$
+$\begin{split}&s_{1}:=\mathbb{E}_{p}[g_{1}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{1}}(x)\cdot g_{1}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\alpha a_{0}+\gamma(a_{1}-a_{0}),\\ &s_{0}:=\mathbb{E}_{p}[g_{0}]=\sum\limits_{x\in\{0,1\}^{n-1}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)\\ &\quad\ =\sum\limits_{x\in A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)+\sum\limits_{x\in A_{1}\setminus A_{0}}\mathbbm{1}_{A_{0}}(x)\cdot g_{0}(x)\mu_{p}^{n-1}(x)=\beta a_{0},\\ &\|g_{1}-g_{0}\|_{2,\mu_{p}}^{2}=\underset{x^{\prime}\sim\mu_{p}^{n-1}}{\mathbb{E}}[(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}]\\ &\quad\ =\sum\limits_{x^{\prime}\in A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})+\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}(g_{1}(x^{\prime})-g_{0}(x^{\prime}))^{2}\mu_{p}^{n-1}(x^{\prime})\\ &\quad\ =(\alpha-\beta)^{2}\cdot a_{0}+(a_{1}-a_{0})\cdot\gamma^{2}.\end{split}$
 Then
@@
 Let $f(t):=\frac{\log_{p}t}{t}$ with $t\in[0,1]$. Then our goal is to prove that
-$ps_{1}^{2}f(a_{1})+(1-p)s_{0}^{2}f(a_{0})+pa_{0}(\alpha-\beta)^{2}+p(a_{1}-a_{0})\gamma^{2}\geq(ps_{1}+(1-p)s_{0})^{2}f(pa_{1}+(1-p)a_{0})).$ (18)
+$\begin{split}ps_{1}^{2}f(a_{1})+(1-p)s_{0}^{2}f(a_{0})&+pa_{0}(\alpha-\beta)^{2}+p(a_{1}-a_{0})\gamma^{2}\\ &\geq(ps_{1}+(1-p)s_{0})^{2}f(pa_{1}+(1-p)a_{0})).\end{split}$ (33)
 Expressing $\beta$ and $\gamma$ as functions of $s_{i},a_{i}$ and of $\alpha$, LHS of the above is a quadratic in $\alpha$ with coefficients depending on $s_{i}$ and $a_{i}$: LHS $A\alpha^{2}+B\alpha+C$, where
@@
 It remains to verify the inequality $4A^{\prime}C^{\prime}\geq B^{\prime 2}$, which reduces to a two-point inequality:
-$pf(a_{1})+(1-p)f(a_{0})+(1-p)a_{1}f(a_{0})f(a_{1})\geq(p(1-p)a_{1}f(a_{0})+(1-p)^{2}a_{1}f(a_{1})+1)f(pa_{1}+(1-p)a_{0}).$ (20)
+$\begin{split}&pf(a_{1})+(1-p)f(a_{0})+(1-p)a_{1}f(a_{0})f(a_{1})\\ &\geq(p(1-p)a_{1}f(a_{0})+(1-p)^{2}a_{1}f(a_{1})+1)f(pa_{1}+(1-p)a_{0}).\end{split}$ (35)
 3.3 Two-point Inequality Part
-Renaming the variables $x=a_{0}$ and $y=a_{1}$, and recalling the constraints on $a_{0}$ and $a_{1}$, we need to prove (20) for $0<x<y\leq 1$.
-$pf(y)+(1-p)f(x)+(1-p)yf(x)f(y)\geq(p(1-p)yf(x)+(1-p)^{2}yf(y)+1)f(py+(1-p)x)$ (21)
+Renaming the variables $x=a_{0}$ and $y=a_{1}$, and recalling the constraints on $a_{0}$ and $a_{1}$, we need to prove Eq. 35 for $0<x<y\leq 1$.
+$\begin{split}&p\cdot f(y)+(1-p)\cdot f(x)+(1-p)\cdot yf(x)f(y)\\ &\geq(p(1-p)yf(x)+(1-p)^{2}yf(y)+1)\cdot f(py+(1-p)x).\end{split}$ (36)
 Rearranging, this is easily seen to be equivalent to
-$\Delta_{p}(x,y)\geq\frac{p(1-p)^{2}y(f(x)-f(y))^{2}}{p(1-p)yf(x)+(1-p)^{2}yf(y)+1}$ (22)
+$\Delta_{p}(x,y)\geq\frac{p(1-p)^{2}y(f(x)-f(y))^{2}}{p(1-p)yf(x)+(1-p)^{2}yf(y)+1},$ (37)
 where
-$\Delta_{p}(x,y):=pf(y)+(1-p)f(x)-f(py+(1-p)x)$
+$\Delta_{p}(x,y):=pf(y)+(1-p)f(x)-f(py+(1-p)x).$
 Note that $\Delta_{p}(x,y)\geq 0$ since $f(t):=\frac{\log_{p}t}{t}$ is convex.
 We now substitute $x=ty$ in $\Delta_{p}(x,y)$, with $0<t<1$, and expand it using Lemma 3.3. We have
-$\begin{split}\Delta_{p}(x,y)=\Delta_{p}(ty,y)&=pf(y)+(1-p)f(t\cdot y)-f((p+(1-p)t)\cdot y)\\ &=pf(y)+\frac{1}{t}\cdot(1-p)f(y)+\frac{1}{y}\cdot(1-p)f(t)-\frac{f(y)}{p+(1-p)t}-\frac{f(p+(1-p)t)}{y}\\ &=\frac{p(1-p)(t-1)^{2}}{t(p+(1-p)t)}\cdot f(y)+\frac{1}{y}\cdot\left(\frac{(1-p)\log_{p}t}{t}-\frac{\log_{p}(p+(1-p)t)}{p+(1-p)t}\right)\end{split}$
-As to the right-hand side of (22), we have
+$\begin{split}&\Delta_{p}(x,y)=\Delta_{p}(ty,y)=pf(y)+(1-p)f(t\cdot y)-f((p+(1-p)t)\cdot y)\\ &=pf(y)+\frac{1}{t}\cdot(1-p)f(y)+\frac{1}{y}\cdot(1-p)f(t)-\frac{f(y)}{p+(1-p)t}-\frac{f(p+(1-p)t)}{y}\\ &=\frac{p(1-p)(t-1)^{2}}{t(p+(1-p)t)}\cdot f(y)+\frac{1}{y}\cdot\left(\frac{(1-p)\log_{p}t}{t}-\frac{\log_{p}(p+(1-p)t)}{p+(1-p)t}\right).\end{split}$
+As to the right-hand side of Eq. 37, we have
 $\begin{split}{\rm RHS}(x,y)&={\rm RHS}(ty,y)=\frac{p(1-p)^{2}y[(\frac{1}{t}-1)f(y)+\frac{1}{y}\cdot\frac{\log_{p}t}{t}]^{2}}{[\frac{p(1-p)}{t}+(1-p)^{2}]yf(y)+p(1-p)\cdot\frac{\log_{p}t}{t}+1}\\ &=\frac{y[(\frac{1}{t}-1)f(y)+\frac{1}{y}\cdot\frac{\log_{p}t}{t}]^{2}}{[\frac{1}{t(1-p)}+\frac{1}{p}]yf(y)+\frac{\log_{p}t}{(1-p)t}+\frac{1}{p(1-p)^{2}}}.\end{split}$
@@
 where
-$z=yf(y)=\log_{p}(y)>0\ \text{as $0<y<1$}$
+$z=yf(y)=\log_{p}(y)>0\ \text{as $0<y<1$},$
 and $A,B,\dots,F$ depend only on $t$ and $p$. Specifically,
@@
 Simplifying and rearranging terms, this inequality reduces to show that
-$G(t,p):=(1-p)(p+(1-p)t)\log_{p}t-p(1-p)\log_{p}(t)\cdot\log_{p}(p+(1-p)t)-t\cdot\log_{p}(p+(1-p)t)\geq 0.$
+$\begin{split}&G(t,p):=(1-p)(p+(1-p)t)\log_{p}t\\ &-p(1-p)\log_{p}(t)\cdot\log_{p}(p+(1-p)t)-t\cdot\log_{p}(p+(1-p)t)\geq 0.\end{split}$
 Clearly, $G(0,p)=G(1,p)=0$.
@@
 We begin by computing the derivative:
-$\begin{split}\frac{\partial}{\partial t}G(t,p)&=\frac{1}{\ln(p)}\left(\frac{p(1-p)}{t}+(1-p)^{2}\ln(t)+(1-p)^{2}-\frac{p(1-p)}{\ln(p)}\cdot\frac{\ln(p+(1-p)t)}{t}\right)\\ &-\frac{1}{\ln(p)}\left(\frac{p(1-p)^{2}}{\ln(p)}\cdot\frac{\ln(t)}{p+(1-p)t}+\ln(p+(1-p)t)+\frac{(1-p)t}{p+(1-p)t}\right).\end{split}$
+$\begin{split}\frac{\partial}{\partial t}G(t,p)&=\frac{1}{\ln(p)}\left(\frac{p(1-p)}{t}+(1-p)^{2}\ln(t)+(1-p)^{2}\right)\\ &-\frac{1}{\ln(p)}\left(\frac{p(1-p)}{\ln(p)}\cdot\frac{\ln(p+(1-p)t)}{t}\right)\\ &-\frac{1}{\ln(p)}\left(\frac{p(1-p)^{2}}{\ln(p)}\cdot\frac{\ln(t)}{p+(1-p)t}+\ln(p+(1-p)t)+\frac{(1-p)t}{p+(1-p)t}\right).\end{split}$
 Define $H(t,p):=t(p+(1-p)t)\cdot\frac{\partial}{\partial t}G(t,p)$. To analyze its sign, we consider successive derivatives of $H$.
-$\begin{split}&\frac{\partial}{\partial t}H(t,p)=\frac{1}{\ln(p)^{2}}\left[-(p-1)^{2}p\left(2+\ln(t)+\ln(p+t-pt)\right)\right]\\ &+\frac{1}{\ln(p)}\left[(-2t+p(2t-1)\ln(p+t-pt)\right]\\ &-\frac{1}{\ln(p)}\left[3(p-1)p(1+p(t-1)-2t)+(p-1)^{2}(-2t+p(2t-1))\ln(t)\right];\\ &\frac{\partial^{2}}{\partial t^{2}}H(t,p)=\frac{1}{\ln(p)^{2}}\left(-(p-1)^{2}p\left(\frac{1}{t}+\frac{1-p}{p+t-pt}\right)\right)\\ &+\frac{1}{\ln(p)}\left(-3(-2+p)(p-1)p-\frac{(p-1)^{2}(-2t+p(2t-1))}{t}+\frac{(1-p)(-2t+p(2t-1))}{p+t-pt}\right)\\ &+\frac{1}{\ln(p)}\left(-2(p-1)^{3}\ln(t)+2(p-1)\ln(p+t-pt)\right);\\ &\frac{\partial^{3}}{\partial t^{3}}H(t,p)=-\frac{(p-1)^{2}p}{t^{2}(p+t-pt)^{2}\ln(p)^{2}}\cdot h(t,p),\end{split}$
+$\begin{split}&\frac{\partial}{\partial t}H(t,p)=\frac{1}{\ln(p)^{2}}\left[-(p-1)^{2}p\left(2+\ln(t)+\ln(p+t-pt)\right)\right]\\ &+\frac{1}{\ln(p)}\left[(-2t+p(2t-1)\ln(p+t-pt)\right]\\ &-\frac{1}{\ln(p)}\left[3(p-1)p(1+p(t-1)-2t)+(p-1)^{2}(-2t+p(2t-1))\ln(t)\right];\\ &\frac{\partial^{2}}{\partial t^{2}}H(t,p)=\frac{1}{\ln(p)^{2}}\left(-(p-1)^{2}p\left(\frac{1}{t}+\frac{1-p}{p+t-pt}\right)\right)\\ &+\frac{1}{\ln(p)}\left(-3(-2+p)(p-1)p-\frac{(p-1)^{2}(-2t+p(2t-1))}{t}\right)\\ &+\frac{1}{\ln(p)}\left(\frac{(1-p)(-2t+p(2t-1))}{p+t-pt}-2(p-1)^{3}\ln(t)+2(p-1)\ln(p+t-pt)\right);\\ &\frac{\partial^{3}}{\partial t^{3}}H(t,p)=-\frac{(p-1)^{2}p}{t^{2}(p+t-pt)^{2}\ln(p)^{2}}\cdot h(t,p),\end{split}$
 where $h(t,p):=-2t^{2}+2pt(-1+2t)+p^{2}(-1+2t-2t^{2})+(-6p(-1+t)t^{2}+4t^{3}+p^{2}(t-1)^{2}(1+2t))\ln(p)$.
@@
 4 Proof of Theorem 1.7
-3.1 Preliminaries
-In this subsection, we outline some notation and terminology, and present some simple lemmas for inductive proof of Theorem 1.8.
-Given $A\subset\{0,1\}^{n}$, we write $x=(x^{\prime},x_{n})\in\{0,1\}^{n}:=V^{n}$. Let
-$V_{j}=\{x\in\{0,1\}^{n}:x_{n}=j\},\ j=0,1$
-and
-$A_{0}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},0)\in A\},\quad A_{1}=\{x^{\prime}\in\{0,1\}^{n-1}:(x^{\prime},1)\in A\}.$
-Let $a_{0}:=\mu^{n-1}(A_{0}),a_{1}:=\mu^{n-1}(A_{1})$, and note that
-$\begin{split}a:=\mu^{n}(A)=\frac{1}{2}\mu^{n-1}(A_{1})+\frac{1}{2}\mu^{n-1}(A_{0})=\frac{a_{0}+a_{1}}{2}.\end{split}$ (23)
-Define $f_{0}:A_{0}\to\mathbb{R},\ f_{1}:A_{1}\to\mathbb{R}$ by
-$f_{0}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},0)\quad\text{and}\quad f_{1}(x_{1},\dots,x_{n-1}):=f(x_{1},\dots,x_{n-1},1).$
-By expanding the Dirichlet form with respect to the conditional expectation on the $n$th variable, we can readily express the Dirichlet form of the original function in terms of those of its restricted versions.
-Lemma 3.1.
-For any $f:A\to\mathbb{R}$, we have
-$\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$ (24)
-Proof.
-We use $x^{\prime}$ to denote the first (n1)(n-1)-coordinates of $x$. Note that
-$\begin{split}\mathcal{E}_{A}(f)&=\frac{1}{4}\cdot\sum\limits_{i=1}^{n-1}\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus i}))^{2}\cdot\mathbbm{1}_{\{x^{\oplus i}\in A\}}\right]+\frac{1}{4}\cdot\underset{x\sim A}{\mathbb{E}}\left[(f(x)-f(x^{\oplus n}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1-x_{n})\in A\}}\right]\\ &=\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},1)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},0)\in A\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime},1)\in A\}}\right]\\ &=\frac{a_{1}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{1}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{1}\}}\right]\\ &+\frac{a_{0}}{2a}\cdot\frac{1}{4}\sum\limits_{i=1}^{n-1}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{0}(x^{\prime\oplus i}))^{2}\cdot\mathbbm{1}_{\{(x^{\prime\oplus i}\in A_{0}\}}\right]\\ &+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right]+\frac{1}{4}\cdot\frac{|A_{0}|}{|A|}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right],\end{split}$
-where in the second equality, we observe that
-$\begin{split}\frac{|A_{1}|}{|A|}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}&\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\right]=\frac{1}{|A|}\cdot\sum\limits_{x^{\prime}\in A_{1}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\cdot\mathbbm{1}_{x^{\prime}\in A_{0}}\\ &=\frac{1}{|A|}\cdot\sum\limits_{x^{\prime}\in A_{0}}(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}=\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{1}(x^{\prime})-f_{0}(x^{\prime}))^{2}\right].\end{split}$
-Similarly, by conditioning on the $n$th coordinate, the variance of the original function can be expressed in terms of the variances of its restricted functions.
-Lemma 3.2.
-For any $f:A\to\mathbb{R}$, we have
-${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (25)
-Proof.
-Note that
-$\begin{split}{\rm Var}_{A}[f]=\underset{x\sim A}{\mathbb{E}}[|f-\underset{x\sim A}{\mathbb{E}}f|^{2}]=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\end{split}$
-Let $\mu=\underset{x\sim A}{\mathbb{E}}[f(x)],\mu_{1}=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]$ and $\mu_{0}=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]$. Then
-$\mu=\frac{1}{|A|}\sum_{x\in A}f(x)=\frac{|A_{1}|}{|A|}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]+\frac{|A_{0}|}{|A|}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]=\frac{a_{1}}{2a}\mu_{1}+\frac{a_{0}}{2a}\mu_{0}.$ (26)
-Thus by the formula $\mathbb{E}[(f-b)^{2}]={\rm Var}[f]+(\mathbb{E}[f]-b)^{2}$, we have that
-$\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]={\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2},$
-and
-$\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]={\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}^{2}}{4a^{2}}(\mu_{1}-\mu_{0})^{2}.$
-Thus we have
-$\begin{split}{\rm Var}_{A}[f]&=\frac{a_{1}}{2a}\cdot\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x\sim A}{\mathbb{E}}f|^{2}]+\frac{a_{0}}{2a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{0}-\underset{x\sim A}{\mathbb{E}}f|^{2}]\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{1}a_{0}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}a_{1}^{2}}{8a^{3}}\cdot(\mu_{1}-\mu_{0})^{2}\\ &=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\cdot(\mu_{1}-\mu_{0})^{2}\end{split}$
-3.2 Induction part
-Our goal is to prove that for every function $f:A\to\mathbb{R}$,
-$\mathcal{E}_{A}(f)\geq c\cdot(1-\sqrt{1-a})\cdot{\rm Var}_{A}[f],$ (27)
-where $c>0$ is a universal constant (we may take $c=0.5$).
-It is easy to verify that the theorem holds for $n=1$.
-Let $n\geq 2$, and suppose the statement of the theorem holds when $n$ is replaced by n1n-1.
-Applying Lemma 3.1 to split the Dirichlet form of the original function into two restricted functions:
-$\mathcal{E}_{A}(f,f)=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right].$ (28)
-By inductive assumption, we have
-$\begin{split}&\mathcal{E}_{A_{1}}(f_{1})\geq c\cdot(1-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}],\\ &\mathcal{E}_{A_{0}}(f_{0})\geq c\cdot(1-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}].\end{split}$ (29)
-Thus our goal is to show that
-$\begin{split}\mathcal{E}_{A}(f)&=\frac{a_{1}}{2a}\cdot\mathcal{E}_{A_{1}}(f_{1})+\frac{a_{0}}{2a}\cdot\mathcal{E}_{A_{0}}(f_{0})+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\geq\frac{a_{1}}{2a}(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}(c-c\sqrt{1-a}){\rm Var}_{A}[f].\end{split}$ (30)
-We also use Lemma 3.2 to split the variance of the original function to two restricted functions:
-${\rm Var}_{A}[f]=\frac{a_{1}}{2a}\cdot{\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}\left(\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})]\right)^{2}.$ (31)
-Thus it suffices to show that
-$\begin{split}&\frac{a_{1}}{2a}\cdot(c-c\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}\cdot(c-c\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{4a}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &\underset{?}{\geq}\frac{a_{1}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{1}}[f_{1}]+\frac{a_{0}}{2a}(c-c\sqrt{1-a}){\rm Var}_{A_{0}}[f_{0}]+\frac{a_{1}a_{0}}{(a_{0}+a_{1})^{2}}(c-c\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (32)
-It is equivalent to show that
-$\begin{split}&ca_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}){\rm Var}_{A_{1}}[f_{1}]+ca_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}){\rm Var}_{A_{0}}[f_{0}]\\ &+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\underset{?}{\geq}c\frac{a_{1}a_{0}}{a_{0}+a_{1}}(1-\sqrt{1-a})(\mu_{1}-\mu_{0})^{2}.\end{split}$ (33)
-Lemma 3.3.
-We may assume both $f_{0},f_{1}$ to be constant on $A_{0}$ and on $A_{1}\setminus A_{0}$.
-Proof.
-Let $\mu_{1}:=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}(x^{\prime})]$, $\alpha:=\mathbb{E}_{x^{\prime}\sim A_{0}}[f_{1}]$, $\beta:=\mathbb{E}_{x^{\prime}\sim A_{1}\setminus A_{0}}[f_{1}]$ and $\gamma:=\mathbb{E}_{x^{\prime}\sim A_{0}}[f_{0}]$.
-Note that
-$\begin{split}\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[|f_{1}-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}f_{1}|^{2}]&=\frac{|A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}+\frac{|A_{1}\setminus A_{0}|}{|A_{1}|}\cdot\frac{1}{|A_{1}\setminus A_{0}|}\sum\limits_{x^{\prime}\in A_{1}\setminus A_{0}}|f_{1}(x^{\prime})-\mu_{1}|^{2}\\ &=\frac{a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}],\end{split}$ (34)
-and
-$\begin{split}&\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}(x^{\prime})-\mu_{0}-(f_{1}(x^{\prime})-\mu_{1})+\mu_{0}-\mu_{1}|^{2}\right]\\ &=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{0}-\mu_{0}|^{2}\right]+\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[|f_{1}-\mu_{1}|^{2}\right]+2\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &+(\mu_{0}-\mu_{1})^{2}+2(\mu_{0}-\mu_{1})(\mu_{1}-\alpha).\end{split}$ (35)
-For the sake of simplicity, we denote $C_{1}=\frac{1}{2}\cdot a_{1}\cdot(\sqrt{1-a}-\sqrt{1-a_{1}})$ and $C_{0}=\frac{1}{2}\cdot a_{0}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}})$. Then
-$\begin{split}&C_{1}\cdot{\rm Var}_{A_{1}}[f_{1}]+C_{0}\cdot{\rm Var}_{A_{0}}[f_{0}]+\frac{a_{0}}{2}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}(x^{\prime})-f_{1}(x^{\prime}))^{2}\right]\\ &=\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+C_{1}\frac{a_{1}-a_{0}}{a_{1}}\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]\\ &+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]+\frac{a_{0}}{2}(\mu_{0}-\mu_{1})^{2}+a_{0}(\mu_{0}-\mu_{1})(\mu_{1}-\alpha)\end{split}$ (36)
-Now we focus on the term
-$\begin{split}&\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]+\left(C_{0}+\frac{a_{0}}{2}\right)\cdot{\rm Var}_{A_{0}}[f_{0}]+a_{0}\cdot\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}\left[(f_{0}-\mu_{0})(\mu_{1}-f_{1})\right]\\ &=\frac{1}{|A_{0}|}\sum\limits_{x^{\prime}\in A_{0}}\left(C_{1}\frac{a_{0}}{a_{1}}+\frac{a_{0}}{2}\right)|\mu_{1}-f_{1}(x^{\prime})|^{2}+\left(C_{0}+\frac{a_{0}}{2}\right)|f_{0}(x^{\prime})-\mu_{0}|^{2}\\ &+a_{0}(f_{0}(x^{\prime})-\mu_{0})(\mu_{1}-f_{1}(x^{\prime}))\\ &=\frac{1}{|A_{0}|}\sum_{x^{\prime}\in A_{0}}\vec{b}_{x^{\prime}}^{\top}G_{x^{\prime}}\vec{b}_{x^{\prime}},\end{split}$ (37)
-where
-$\vec{b}_{x^{\prime}}=\begin{bmatrix}f_{0}(x^{\prime})-\mu_{0}\\ \mu_{1}-f_{1}(x^{\prime})\end{bmatrix},\\ G_{x^{\prime}}=\begin{bmatrix}a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}&\frac{a_{0}}{2}\\ \frac{a_{0}}{2}&a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}\end{bmatrix}$.\\
-We claim that the matrix $G_{x^{\prime}}$ is positive semidefinite.
-Note that $G_{x^{\prime}}$ is a symmetric matrix, then we shall use the basic fact that a symmetric matrix $G=\begin{bmatrix}b&c\\ c&d\end{bmatrix}\in\mathbb{R}^{2\times 2}$ is positive semidefinite if and only if $b\geq 0,d\geq 0$ and $bd-c^{2}\geq 0$.
-It is easy to see that $a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{0}}+1}{2}\geq 0$ and $a_{0}\cdot\frac{\sqrt{1-a}-\sqrt{1-a_{1}}+1}{2}\geq 0$ as $a_{0}\leq a\leq a_{1}$. Thus it suffices to verify
-$\frac{a_{0}^{2}}{4}\cdot(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-\frac{a_{0}^{2}}{4}\geq 0.$ (38)
-Let $s:=\sqrt{1-a_{0}}$ and $t:=\sqrt{1-a_{1}}$, then $0\leq t\leq s\leq 1$, $1-a=(s^{2}+t^{2})/2$ and
-$\begin{split}&(\sqrt{1-a}-\sqrt{1-a_{0}}+1)\cdot(\sqrt{1-a}-\sqrt{1-a_{1}}+1)-1\\ &=\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-s+1\right)\cdot\left(\sqrt{\frac{s^{2}+t^{2}}{2}}-t+1\right)-1\\ &=-\frac{1}{2}\cdot(s+t-2)\cdot(\sqrt{2}\sqrt{s^{2}+t^{2}}-s-t)\geq 0,\end{split}$ (39)
-where the final one is obviously nonnegative as $s+t-2\leq 0$ and $2(s^{2}+t^{2})\geq(s+t)^{2}$.
-By the second-order condition of convexity, we know that for any $x^{\prime}\in A_{0}$, the quadratic function $G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))$ is a convex function.
-Thus by Jensen’s inequality, we have
-$\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[G(f_{0}(x^{\prime})-\mu_{0},\mu_{1}-f_{1}(x^{\prime}))]\geq G\left(\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}(x^{\prime})-\mu_{0}],\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[\mu_{1}-f_{1}(x^{\prime})]\right)=G(0,\mu_{1}-\alpha),$
-and
-$\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[|f_{1}-\mu_{1}|^{2}]\geq\left(\underset{x^{\prime}\sim A_{1}\setminus A_{0}}{\mathbb{E}}[f_{1}-\mu_{1}]\right)^{2}=(\beta-\mu_{1})^{2}.$
-Hence replacing $f_{j}$ with their averages on the corresponding subsets can only decrease the left-hand side and does not change the right-hand side in the second inequality of (33). ∎
-3.3 Five-point Inequality Part
-Now let $\alpha$ be the value of $f_{1}$ on $A_{0}$ and $\beta$ be the value of $f_{1}$ on $A_{1}\setminus A_{0}$. Let $\gamma$ be the value of $f_{0}$ on $A_{0}$. Then we compute
-$\begin{split}&{\rm Var}_{A_{1}}[f_{1}]=\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}^{2}]-\underset{x^{\prime}\sim A_{1}}{\mathbb{E}}[f_{1}]^{2}\\ &=\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta^{2})-\left(\frac{1}{2^{n-1}\cdot a_{1}}(a_{0}\cdot 2^{n-1}\cdot\alpha+(a_{1}-a_{0})\cdot 2^{n-1}\cdot\beta)\right)^{2}\\ &=\frac{1}{a_{1}}(a_{0}\cdot\alpha^{2}+(a_{1}-a_{0})\cdot\beta^{2})-\frac{1}{a_{1}^{2}}(a_{0}\cdot\alpha+(a_{1}-a_{0})\cdot\beta)^{2}\\ &=\frac{a_{1}a_{0}-a_{0}^{2}}{a_{1}^{2}}\cdot(\alpha-\beta)^{2},\end{split}$ (40)
-and
-${\rm Var}_{A_{0}}[f_{0}]=\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}^{2}]-\underset{x^{\prime}\sim A_{0}}{\mathbb{E}}[f_{0}]^{2}=\frac{a_{0}\cdot 2^{n-1}\cdot\gamma^{2}}{2^{n-1}\cdot a_{0}}-\left(\frac{a_{0}\cdot 2^{n-1}\cdot\gamma}{2^{n-1}\cdot a_{0}}\right)^{2}=0.$ (41)
-Thus to verify (33), it suffices to prove the following five-point inequality with $c=\frac{1}{2}$.
-$\begin{split}&c(\sqrt{1-a}-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})\cdot(\alpha-\beta)^{2}+\frac{a_{1}}{2}\cdot(\gamma-\alpha)^{2}\\ &\geq\frac{c}{a_{1}+a_{0}}\cdot(1-\sqrt{1-a})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}.\end{split}$ (42)
-Note that
-$\sqrt{1-a}-\sqrt{1-a_{1}}=\frac{1-a-(1-a_{1})}{\sqrt{1-a}+\sqrt{1-a_{1}}}=\frac{a_{1}-\frac{a_{1}+a_{0}}{2}}{\sqrt{1-a}+\sqrt{1-a_{1}}}=\frac{a_{1}-a_{0}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})},$ (43)
-and
-$1-\sqrt{1-a}=\frac{1-(1-a)}{1+\sqrt{1-a}}=\frac{a_{1}+a_{0}}{2(1+\sqrt{1-a})}.$ (44)
-Then it suffice to prove that
-$\begin{split}&\frac{c\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}}{2(\sqrt{1-a}+\sqrt{1-a_{1}})}+\frac{a_{1}}{2}\cdot(\alpha-\gamma)^{2}\geq\frac{c}{2(1+\sqrt{1-a})}\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\\ &\Longleftrightarrow c\cdot(1+\sqrt{1-a})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot[a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)]^{2}\geq 0.\end{split}$ (45)
-Note that
-$a_{1}\cdot(\beta-\gamma)+a_{0}(\alpha-\beta)=(a_{0}-a_{1})\cdot(\alpha-\beta)+a_{1}\cdot(\alpha-\gamma)$
-After some simple calculations, the inequality to be verified becomes
-$\begin{split}&c(1-\sqrt{1-a_{1}})\cdot(a_{1}-a_{0})^{2}\cdot(\alpha-\beta)^{2}+a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot(\alpha-\gamma)^{2}\\ &-2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\cdot(a_{0}-a_{1})\cdot(\alpha-\beta)\cdot(\alpha-\gamma)\geq 0.\end{split}$ (46)
-If $\alpha=\gamma$, the inequality is trivially satisfied.
-Thus, assume $\alpha\neq\gamma$ and set $T:=\frac{\alpha-\beta}{\alpha-\gamma}\cdot(a_{1}-a_{0})$.
-Then the above inequality can be rewritten as a quadratic in $T$:
-$\begin{split}&c\cdot(1-\sqrt{1-a_{1}})\cdot T^{2}+2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot T\\ &a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\geq 0\end{split}$ (47)
-We may rewrite the left-hand side as a quadratic polynomial in $T$ ($AT^{2}+BT+C$), whose coefficients depend only on $a_{1},a_{0}$:
-$\begin{split}&A:=c\cdot(1-\sqrt{1-a_{1}})\geq 0\\ &B:=2\cdot c\cdot(\sqrt{1-a}+\sqrt{1-a_{1}})\cdot a_{1}\\ &C:=a_{1}\cdot(1+\sqrt{1-a}-c\cdot a_{1})\cdot(\sqrt{1-a}+\sqrt{1-a_{1}}).\end{split}$
-To ensure the quadratic is nonnegative for all $T$, it suffices to check that its discriminant satisfies $\Delta=B^{2}-4AC\leq 0$ with $c=0.5$, which will conclude the proof.
-Lemma 3.4.
-$B^{2}-4AC\leq 0$.
-Proof.
-Denote $x=a_{0},y=a_{1}$. Simplifying and rearranging, we need to prove that for any $0\leq x\leq y\leq 1$, we have
-$\begin{split}&y\cdot\left(\sqrt{1-\frac{x+y}{2}}+\sqrt{1-y}\right)-2(1-\sqrt{1-y})\cdot\left(1+\sqrt{1-\frac{x+y}{2}}-\frac{y}{2}\right)\leq 0\\ &\Leftrightarrow\left(\sqrt{1-\frac{x+y}{2}}+1\right)\cdot(y-2+2\sqrt{1-y})\leq 0.\end{split}$ (48)
-It suffices to show that $(y-2+2\sqrt{1-y})\leq 0$. It is trivially true as
-$\begin{split}&y-2+2\sqrt{1-y}\leq 0\\ &\Leftrightarrow 2\sqrt{1-y}\leq 2-y\\ &\Leftrightarrow y^{2}\geq 0.\end{split}$ (49)
+We now show how to derive Theorem 1.7 from the $p$-biased discrete functional inequality Eq. 8 established in Theorem 1.6.
+For $g:\{0,1\}^{n}\to\mathbb{R}$, recall the $p$-biased Dirichlet form
+$\mathcal{E}_{p}(g,g):=\sum_{x}\mu_{p}(x)\sum_{y\sim x}\left(g(x)-g(y)\right)^{2}.$
+Let $L_{p}:=D-W$ be the (external) weighted Laplacian on $A$. Then the standard edge-vertex double counting gives
+$\mathcal{E}_{p}(g,g)=\sum_{\{x,y\}\subseteq A}\omega_{p}(x,y)\bigl(g(x)-g(y)\bigr)^{2}=\vec{g}^{\top}L_{p}\vec{g}.$ (38)
+Indeed, grouping the two directed contributions $(x,y)$ and $(y,x)$ across each undirected edge $\{x,y\}$ yields
+$\begin{split}\mu_{p}(x)(g(x)-g(y))^{2}+\mu_{p}(y)(g(x)-g(y))^{2}&=(\mu_{p}(x)+\mu_{p}(y))(g(x)-g(y))^{2}\\ &=\omega_{p}(x,y)(g(x)-g(y))^{2};\end{split}$
+expanding the rightmost equality is the usual identity $\sum_{\{x,y\}}\omega_{p}(x,y)(g(x)-g(y))^{2}=\vec{g}^{\top}(D-W)\vec{g}$.
+We may assume without loss of generality that $g\geq 0$.
+Indeed, since $||g(x)|-|g(y)||\leq|g(x)-g(y)|$, replacing $g$ by $|g|$ can only decrease the Dirichlet energy, while $\mathbb{E}_{\mu_{p}}[|g|]\geq|\mathbb{E}_{\mu_{p}}[g]|$ ensures that the right-hand side of Eq.
+8 does not decrease.
+Therefore, Theorem 1.6 gives
+$p\cdot\mathcal{E}_{p}(g,g)\geq\frac{(\mathbb{E}_{\mu_{p}}g)^{2}}{\mu_{p}(A)}\cdot\log_{p}(\mu_{p}(A)).$ (39)
+Let $\vec{\mu}=(\mu_{p}(x))_{x\in A}$, so that $\mathbb{E}_{\mu_{p}}[g]=\left\langle\vec{\mu},\vec{g}\right\rangle$. Inserting (38) into (39) gives
+$\vec{g}^{\top}L_{p}\vec{g}\ \geq\rho\cdot(\vec{\mu}^{\top}\vec{g})^{2},\quad\rho:=\frac{\log_{p}\left(\mu_{p}(A)\right)}{p\mu_{p}(A)}.$
+Equivalently,
+$L_{p}\succeq\rho\cdot\vec{\mu}\vec{\mu}^{\top}\Rightarrow\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\leq\frac{1}{\rho}=\frac{p\mu_{p}(A)}{\log_{p}\big(\mu_{p}(A)\big)}.$ (40)
+Since $L_{p}=D(I-P_{A})$, we have
+$L_{p}^{-1}=(I-P_{A})^{-1}D^{-1}=\sum_{k\geq 0}P_{A}^{k}D^{-1}.$ (41)
+Let $\vec{\nu}$ be the initial distribution $\nu(x)=\mu_{p}(x)/\mu_{p}(A)$ on $A$ and set $\vec{\phi}:=D^{-1}\vec{\mu}$, i.e.
+$\phi(x)=\mu_{p}(x)/d_{p}(x)$.
+Let $(X_{t})_{t\geq 0}$ be the Markov chain with transition $P_{A}$, killed upon exiting $A$, and let $\tau:=\min\{t\geq 1:X_{t}\notin A\}$.
+By the tower property and the Markov property, for every $k\geq 0$ and every $f:\{0,1\}^{n}\to\mathbb{R}$ supported on $A$,
+$(P_{A}^{k}f)(x)\ =\ \mathbb{E}_{x}\left[f(X_{k})\mathbbm{1}_{\{\tau>k\}}\right],$ (42)
+where $\mathbb{E}_{x}$ is the expectation taken with respect to a $p$-biased random walk starting at a given point $x$. Applying (42) with $f=\phi$ and averaging over $\vec{\nu}$,
+$\vec{\nu}^{\top}P_{A}^{k}\vec{\phi}=\mathbb{E}_{\nu}\left[\phi(X_{k})\mathbbm{1}_{\{\tau>k\}}\right].$
+Using (41) we obtain
+$\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}=\sum_{k\geq 0}\vec{\mu}^{\top}P_{A}^{k}D^{-1}\vec{\mu}=\mu_{p}(A)\sum_{k\geq 0}\vec{\nu}^{\top}P_{A}^{k}\vec{\phi}=\mu_{p}(A)\sum_{k\geq 0}\mathbb{E}_{\nu}\left[\phi(X_{k})\mathbbm{1}_{\{\tau>k\}}\right].$ (43)
+Combining with Eq. 40 proves
+$\tilde{T}(A,p)=\sum_{k\geq 0}\mathbb{E}_{\nu}\left[\phi(X_{k})\mathbbm{1}_{\{\tau>k\}}\right]=\frac{1}{\mu_{p}(A)}\cdot\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\leq\frac{p}{\log_{p}\big(\mu_{p}(A)\big)},$
+where the equality case is exactly the extremal case of Theorem 1.6, namely that $A$ is an increasing subcube.
+For $x$ with Hamming weight $t=|x|$, a direct calculation gives
+$\phi(x)=\frac{\mu_{p}(x)}{d_{p}(x)}\ =\ \frac{1}{\,n+(n-t)\frac{p}{1-p}+t\frac{1-p}{p}\,}.$ (44)
+Hence for all $x\in A$,
+$\frac{\min\{p,1-p\}}{n}\ \leq\ \phi(x)\ \leq\ \frac{\max\{p,1-p\}}{n}.$ (45)
+From (43) and (45) we get
+$\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\geq\frac{\min\{p,1-p\}}{n}\mu_{p}(A)\sum_{k\geq 0}\mathbb{P}_{\nu}(\tau>k)=\frac{\min\{p,1-p\}}{n}\mu_{p}(A)\mathbb{E}_{\nu}[Y].$ (46)
+Combining (40) and (46) yields
+$\frac{\min\{p,1-p\}}{n}\mu_{p}(A)\mathbb{E}_{\nu}[Y]\leq\vec{\mu}^{\top}L_{p}^{-1}\vec{\mu}\leq\frac{p\mu_{p}(A)}{\log_{p}\big(\mu_{p}(A)\big)},$
+hence
+$T(A,p)=\mathbb{E}_{\nu}[Y]\leq\frac{p}{\min\{p,1-p\}}\cdot\frac{n}{\log_{p}\big(\mu_{p}(A)\big)}.$
+Equality case. Equality in the final bound requires equality in each step where an inequality was used:
+(a) Equality in (40) is equivalent to equality in the functional inequality (39). By the equality characterization of Theorem 1.6, this holds if and only if $A$ is an increasing subcube.
+(b) Equality in (46) forces $\phi(X_{k})$ to be almost surely constant along all surviving paths.
+From (44) one sees that $\phi$ depends linearly on $|x|$ and is constant on AA only when $p=\tfrac{1}{2}$ (in which case $\phi\equiv 1/(2n)$).
+Thus the equality case occurs exactly when $p=\tfrac{1}{2}$ and $A$ is an increasing subcube.
+5 Proof of mean first exit time for increasing subcubes
+In this section, we verify the mean survival time of random walks from increasing subcubes (1).
+For $x\in A$, write $t(x):=\sum_{i=k+1}^{n}x_{i}$ for the number of free $1$’s.
+By permutation symmetry among the free coordinates, the conditional expectation of the exit time only depends on $t(x)$.
+Set
+$m(t):=\mathbb{E}[Y\mid t(X_{0})=t],\ t=0,1,\dots,n-k.$
+From a state with value $t$, exactly three mutually exclusive one-step outcomes can occur:
+  • (exit) flip one of the first $k$ fixed $1$’s to 0 — this exits $A$;
+  • (down) flip a free $1$ to 0, so $t\mapsto t-1$;
+  • (up) flip a free 0 to $1$, so $t\mapsto t+1$.
+Counting the three kinds of neighbors (there are $k$ exit-edges, $t$ down-edges, and nktn-k-t up-edges), we obtain
+$\frac{d_{p}(x)}{\mu_{p}(x)}=\frac{k+t}{p}+\frac{n-k-t}{1-p}.$
+Define
+$D(t):=p(1-p)\frac{d_{p}(x)}{\mu_{p}(x)}=(1-p)(k+t)+p\bigl((n-k)-t\bigr),$
+and hence the three one-step probabilities from level $t$ are
+$e(t)=\frac{k(1-p)}{D(t)},\ d(t)=\frac{t(1-p)}{D(t)},\ u(t)=\frac{(n-k-t)p}{D(t)},\ e(t)+d(t)+u(t)=1.$
+By the Markov property (first-step analysis), conditioning on the three one-step outcomes gives, for every $t\in\{0,\dots,n-k\}$,
+$m(t)=1+d(t)\cdot m(t-1)+u(t)\cdot m(t+1),$ (47)
+with d(0)=u(nk)=0d(0)=u(n-k)=0. Now let us compute the equation
+$D(t)m(t)=D(t)+t(1-p)\cdot m(t-1)+(n-k-t)p\cdot m(t+1).$
+Substituting $m(t)=A+Bt$, expanding both sides and matching coefficients of $t^{2},t^{1}$ and the constant term yields
+$B=\frac{1-2p}{1+(1-p)k},\quad A=1+\frac{p(n-k)}{(1-p)k}(1+B).$
+Under the starting law $\nu=\mu_{p}(\cdot)/\mu_{p}(A)$, the free coordinates are i.i.d. ${\rm Ber}(p)$, so $T_{0}:=t(X_{0})\sim{\rm Bin}(n-k,p)$ and $\mathbb{E}[T_{0}]=p(n-k)$. Hence
+$T(C,p)=\mathbb{E}_{\nu}[Y]=\mathbb{E}[m(T_{0})]=A+B\cdot\mathbb{E}[T_{0}]=1+2p\left(\frac{n}{k}-1\right).$
+This completes the proof.
 6 Conclusion and open problems
-The weak Harper’s edge-isoperimetric inequality (2) has several easy proofs based on induction.
-Once using conditional expectation to expand the total influence and by inductive assumption, it remains to check that the following two-point inequality for $a_{0},a_{1}\in(0,1)$
-$a_{1}\log_{2}\left(\frac{1}{a_{1}}\right)+a_{0}\log_{2}\left(\frac{1}{a_{0}}\right)+|a_{1}-a_{0}|\geq(a_{1}+a_{0})\log_{2}\left(\frac{2}{a_{1}+a_{0}}\right),$
-which is not hard to verify, using, e.g., concavity of the entropy function $H(x)=-x\log x-(1-x)\log(1-x)$ on $[0,1]$.
-It may be more amazing that we can prove the ‘sharp’ Harper’s edge-isoperimetric inequality by using entropy method, see Section 4.4 in [7].
-Since numerous classical results in Boolean function analysis–including the hypercontractivity inequality–can be established via entropy methods, see for example [32], we are very interested in whether other methods, especially analytical methods (or spectral graph theory), can be used to prove tight Harper’s edge-isoperimetric inequality.
-A natural next target beyond our $O(n^{2})$ mixing bound for constant-density increasing sets is the full Ding–Mossel conjecture asking for an $O(n\log n)$ mixing time for the censored walk when $\mu(A)\geq\varepsilon>0$ ([10, Question 1.1]).
-As observed in [17], a route to such an $O(n\log n)$ bound is to establish an appropriate (standard or modified) log-Sobolev inequality for the censored chain on $A$.
-Following the general relations between log-Sobolev and mixing time bound (e.g.,[6, 9]), it is natural to ask what the correct dependence on the density $\mu(A)$ should be in any valid log–Sobolev (or modified log-Sobolev) inequality for increasing sets.
-Our inductive two-point machinery is well suited for quadratic (L2) control: it converts high-dimensional Dirichlet forms and variances into lower-dimensional analogues plus a finite list of analytics inequalities (e.g., two-point inequalities) that one can verify exactly.
+A natural next target beyond our $O(n^{2})$ mixing bound for constant-density increasing sets is the full Ding–Mossel conjecture asking for an $O(n\log n)$ mixing time for the censored walk when $\mu(A)\geq\varepsilon>0$ ([DM2014, Question 1.1]).
+As observed in [fei2025spectral], a route to such an $O(n\log n)$ bound is to establish an appropriate (standard or modified) log-Sobolev inequality for the censored chain on $A$.
+Following the general relations between log-Sobolev and mixing time bound (e.g.,[BT2006logsobolev, DS1996logsobolev]), it is natural to ask what the correct dependence on the density $\mu(A)$ should be in any valid log–Sobolev (or modified log-Sobolev) inequality for increasing sets.
+Our inductive two-point machinery is well suited for quadratic ($L^{2}$) control: it converts high-dimensional Dirichlet forms and variances into lower-dimensional analogues plus a finite list of analytics inequalities (e.g., two-point inequalities) that one can verify exactly.
 However, in log-Sobolev inequalities, the relative entropy is a nonlinear, non-quadratic functional—this introduces challenges when applying the induction method to control it.